### v1

# Moduli spaces of sextic curves with simple singularities and their compactifications

Abstract.
In this paper, we study moduli spaces of sextic curves with simple singularities. Through period maps of K3 surfaces with ADE singularities, we prove such moduli spaces admit algebraic open embeddings into arithmetic quotients of type IV domains. For all cases, we prove the identifications of GIT compactifications and Looijenga compactifications. We also describe Picard lattices in an explicit way for many cases and apply this to study the relation of orbifold structures on two sides of the period map.
1. Introduction
The study of moduli spaces of singular plane curves has a long history dating back to Severi [Sev21], where he demonstrated that families of nodal plane curves of fixed degree form smooth varieties of the expected dimension. In this paper, we focus on sextic curves. A double cover of $\mathbb{P}^{2}$ branched along a smooth sextic curve is a K3 surface of degree two. Thanks to the global Torelli theorem for K3 surfaces, one can identify the moduli of smooth sextic curves with the complement of two irreducible Heegner divisors in an arithmetic quotient of a type IV domain. In his seminal work [Sha80], Shah described stable and semi-stable sextic curves. Based on this, there is an identification between the GIT compactification and the Looijenga compactification [Loo03] of the moduli of sextic curves.
In this paper we study moduli of sextic curves with simple singularities. Previously, the first two authors [YZ23] considered moduli spaces of nodal sextic curves. The main result of [YZ23] is the identification of GIT compactifications with Looijenga compactifications for such moduli spaces. Additionally, the Picard lattices of the K3 surfaces associated to generic nodal sextic curves with given singular type are determined. In this paper, we generalize the main results in [YZ23] to the cases of sextic curves with simple singularities.
Given a root system $R$ with ADE components, the sextic curves with ADE singularities of type $R$ form a subvariety of $\mathbb{P}^{27}=\mathbb{P}\mathrm{Sym}^{6}(\mathbb{C}^{3})^{\vee}$, see [GL96, Prop 2.1]. An irreducible component of this subvariety is called a singular type, denoted by $T$. Greuel–Lossen’s result [GL96, Corollary 5.1(b)] and [GS21, Theorem 2.5.1] implies that each irreducible component is smooth of the expected dimension. Thus, for any given singular type $T$, we can define the moduli space $\mathcal{M}_{T}\coloneqq\mathrm{PGL}(3)\backslash\!\!\backslash\mathcal{V}_{T}$, where $\mathcal{V}_{T}$ denotes the associated irreducible component. All such $R$ are classified by Urabe [Ura88] and Yang [Yan96], where the maximal rank is $19$. By Yang [Yan96, Thm 2.1], the number of $R$ with rank $19$ (respectively, $18$, $17$, $16$) is $519$ (respectively, $987$, $975$, $782$).
Next we formulate our main theorem and highlight the key ingredients for the proof. Fix a singular type $T$ (with root system $R$) and take a sextic curve $Z$ of such type. Let $X$ be the minimal resolution of the double cover of $\mathbb{P}^{2}$ branched along $Z$. Then $X$ is a K3 surface. The Picard lattice $\mathrm{Pic}(X)$ naturally contains a degree two class $H$ (which is the pullback of the hyperplane class of $\mathbb{P}^{2}$) and a root lattice $L$ of type $R$. Let $P$ be the primitive hull of $\langle H\rangle\oplus L$ in $\mathrm{Pic}(X)$. The orthogonal complement $Q$ of $P$ in $\Lambda_{X}\coloneqq H^{2}(X,\mathbb{Z})$ is a lattice with signature $(2,20-\mathrm{rank}P)$, thus defines a type IV domain $\mathbb{D}_{T}\coloneqq\mathbb{D}(Q)$. The period of $P$-polarized K3 surfaces (see also Dolgachev [Dol96]) gives rise to an algebraic morphism
$\mathscr{P}_{T}\colon\mathcal{M}_{T}\to\Gamma_{T}\backslash\mathbb{D}_{T}.$
Here the monodromy group $\Gamma_{T}$ is defined in §3.1. We call $\mathscr{P}_{T}$ an occult period map, following Kudla–Rapoport [KR12], to distinguish with the usual period map.
Let $\mathcal{H}_{T}$ be the union of hyperplanes with normal vectors in $\{r\in H^{\perp}\,|\,r^{2}=-2,r\notin L\}$. Let $\mathcal{H}_{T}^{*}$ be the sub-arrangement of $\mathcal{H}_{T}$ defined by roots with divisibility two. Both $\mathcal{H}_{T}$ and $\mathcal{H}_{T}^{*}$ are $\Gamma_{T}$-invariant. One of the main results of this paper is the following, see Theorem 3.3, Theorem 3.4, Theorem 4.12 and Proposition 6.6.
Theorem  1.1.
The occult period map $\mathscr{P}_{T}\colon\mathcal{M}_{T}\to\Gamma_{T}\backslash\mathbb{D}_{T}$ is an open embedding whose image is $\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$. It extends to an isomorphism
¯$\overline{\mathcal{M}}_{T}\cong\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathcal{H}_{T}^{*}},$
where ¯$\overline{\mathcal{M}}_{T}$ is the normalization for the GIT compactification of $\mathcal{M}_{T}$, and $\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathcal{H}_{T}^{*}}$ denotes the Looijenga compactification of $\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T}^{*})$. Moreover, when $T$ is a nodal singular type, $\mathscr{P}_{T}$ induces an isomorphism of $\mathcal{M}_{T}$ and $P\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$ as orbifolds.
There are two main ingredients in the proof for the injectivity of $\mathscr{P}_{T}$. First, we need to find the correct subgroup $\Gamma_{T}$ of $\mathrm{O}(Q)$. It turns out that $\Gamma_{T}$ is the group of automorphisms of $Q$ that can be extended to automorphisms of the K3 lattice preserving $H$ and certain base of $L$, see §3.1. Under this definition, the group $\Gamma_{T}$ is sufficiently large to make $\mathscr{P}_{T}$ well-defined. Meanwhile, it is not too large so that we can actually prove injectivity of $\mathscr{P}_{T}$. Next, for each $F\in\mathcal{V}_{T}$, we construct an ample class on the K3 surface $X_{F}$ in a uniform way. Combining our characterization of $\Gamma_{T}$ and the construction of the ample class, the injectivity basically follows from the global Torelli theorem for K3 surfaces. Since $\mathcal{M}_{T}$ is of the the expected dimension, the period map is an open embedding. A corollary is that $\mathrm{Pic}(X)$ associated with a generic $Z\in\mathcal{M}_{T}$ is $P$.
Next, we explain the characterization for the image of the period map. Some arguments here are originally due to Urabe [Ura88] and Degtyarev [Deg08], see §3.3. For any $\omega\in\mathbb{D}_{T}-\mathcal{H}_{T}$, there exists a K3 surface $X_{\omega}$ by the surjectivity of the period map of K3 surfaces. We show that $X_{\omega}$ arises from a sextic curve $Z_{\omega}$ of singular type $T$ and $\mathscr{P}_{T}(Z_{\omega})=[\omega]$ in $\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$. The proof follows from the following two steps. First, we construct a nef class $H_{\omega}$ on $X_{\omega}$ and show that its complete linear system induces a surjective morphism $\pi_{\omega}\colon X_{\omega}\to\mathbb{P}^{2}$ of degree two. This morphism contracts exactly the rational curves orthogonal to $H_{\omega}$. Then $\pi_{\omega}$ factors through a finite morphism $p_{\omega}\colon\widehat{X}_{\omega}\to\mathbb{P}^{2}$ which is a double cover branched along a sextic curve $Z_{\omega}$. Next, the automatic semiregularity of a curve in a K3 surface allows us to construct $\pi_{\omega}$ in families. This concludes that $Z_{\omega}$ has the singular type $T$.
For identification between the GIT and Looijenga compactifications, we follow the approach developed in [YZ20, YZ23]. The proof can be visualized in Diagram (1). Here $\mathbb{D}_{1}$ and $\Gamma_{1}$ are the type IV domain and monodromy group associated to smooth sextic curves (namely $T$ is the smooth type), and $\mathcal{H}_{\infty}$ is defined by roots of divisibility two, see §4.1. We have already shown that $\mathscr{P}_{T}$ is an isomorphism between $\mathcal{M}_{T}$ and $\Gamma\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$. There is a natural map $\pi_{T}\colon\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T}^{*})\to\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\infty})$. We claim that $\pi_{T}$ extends to a finite morphism between Looijenga compactifications. The key point is the functorial property of Looijenga compactifications $\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathcal{H}_{T}^{*}}$ and $\overline{\Gamma_{1}\backslash\mathbb{D}_{1}}^{\mathcal{H}_{\infty}}$, see [Loo03] and [YZ20, Theorem A.13]. To fulfill the conditions in [YZ20, Theorem A.13], we need to check two facts. Firstly, the stabilizer of a generic point of $\mathbb{D}_{T}$ in $\mathrm{O}(\Lambda_{X},H)$ is generated by the Weyl group $W(R)$ and $-\mathrm{Id}$. Secondly, we show that $\Gamma_{T}$ is the normalizer of $W(R)$ in $\Gamma_{1}$. Then we conclude that $\mathscr{P}_{T}$ extends to an isomorphism between ¯$\overline{\mathcal{M}}_{T}$ and $\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathcal{H}_{T}^{*}}$ by showing that both vertical maps $j_{T}$ and $\pi_{T}$ are normalizations onto their images.
[…]
The last assertion on orbifold structures is proved based on a lattice-theoretic argument, see Proposition 6.6.
In §5, we present some further analysis of the generic Picard lattice $P$. We first consider the lattice $M$ generated by $H$, $L$ and the strict transforms of irreducible components of the sextic curve under $X\to\mathbb{P}^{2}$. Apparently $P$ is a saturation of $M$, and a natural question is when $P=M$ holds. This is still open in general, and in §5 we prove that $P^{\iota}=M^{\iota}$ for the natural involution $\iota$ on $\Lambda_{X}$, see Proposition 5.9.
Structure of the paper. In §2, we use GIT to characterize the moduli space of sextic curves for a given singular type. Then we review the resolution of ADE singularities and apply it to the case of double plane branched over an ADE sextic curve. In §3, we construct the period map $\mathscr{P}_{T}$ for the moduli space of sextic curves of a given singular type $T$ and prove the injectivity of $\mathscr{P}_{T}$. Moreover, we characterize the image of $\mathscr{P}_{T}$ and conclude that $\mathscr{P}_{T}$ is an open embedding. In §4, we establish a correspondence between the GIT compactification of the moduli space and the Looijenga compactification of the arithmetic quotient. Moreover, we show that the compactifications have a natural functorial property. In §5, we discuss the generic Picard lattice $P$. We analyze the contributions of the irreducible components to $P$ and show that the invariant part $P^{\iota}$ is generated by the hyperplane class, the invariant part of root lattices and the classes of irreducible components. In §6, we establish a criterion for whether the period map preserves the orbifold structures of two sides. We show that $\mathscr{P}_{T}$ preserves the orbifold structures for all nodal singular types $T$. Finally in §7, we apply our results to some interesting cases, including a revisit of the Zariski pairs.
Acknowledgement. The first author is supported by the national key research and development program of China (No. 2022YFA1007100) and NSFC 12201337. The second author is partially supported by NSFC 12301058. We thank Bong Lian, Zhiyu Tian and Zheng Zhang for their interests and helpful discussions.
2. ADE Sextic Curves and K3 Surfaces
2.1. Plane Curves with Simple Singularities
We always use $Z$ to represent a reduced complex plane curve of positive even degree. We denote by $\widehat{X}$ the double cover of $\mathbb{P}^{2}$ branched along $Z$ and let $X$ be the smooth minimal model of $\widehat{X}$. Let $R$ be an irreducible root system of ADE type. Suppose $p\in Z$ is a singular point and the corresponding singularity in $\widehat{X}$ is an ADE singularity of type $R$, then $p\in Z$ is a simple curve singularity of type $R$.
Suppose $Z$ has only simple singularities. A plane curve $Z^{\prime}$ is said to have the same singular type as $Z$ if they are equisingular deformation equivalent. For a general discussion of equisingularity, see [Zar65a, Zar65b, Zar68] and [GLS07]. We have introduced an alternative definition of singular type in §1. The two definitions are equivalent due to [GL96, Proposition 2.1]. Denote by $T$ the singular type of $Z$. Let $\mathbb{P}\mathcal{V}_{T}$ be the space for curves of a fixed singular type $T$, where $\mathcal{V}_{T}$ denotes the cone of polynomials over $\mathbb{P}\mathcal{V}_{T}$. By [GL96, Prop 2.1], $\mathbb{P}\mathcal{V}$ is an irreducible quasi-projective variety.
We define the expected dimension of $\mathbb{P}\mathcal{V}_{T}$. Let $d=\deg(Z)$. Let $x_{1},\cdots,x_{n}$ be the set of singularities of $Z$. Denote by $R_{i}$ the corresponding ADE type of $x_{i}$. The number $\mathrm{rank}(R_{i})$ is also called the Milnor number or the Tjurina number for the corresponding surface singularity in other contexts. The expected dimension of $\mathbb{P}\mathcal{V}_{T}$ is defined to be $\mathrm{edim}(\mathbb{P}\mathcal{V})=\binom{d+2}{2}-1-\sum\limits_{i=1}^{n}\mathrm{rank}(R_{i})$. When equality holds, we simply say that $\mathbb{P}\mathcal{V}_{T}$ is of the expected dimension.
When $d=6$, the space $\mathbb{P}\mathcal{V}_{T}$ is a smooth irreducible quasi-projective variety of the expected dimension, see [GL96, Corollary 5.1(b)].
2.2. Resolution of ADE Singularities
In this section, we review the resolution of surface singularities of ADE types. Each such singularity can be realized as $o=(0,0,0)$ in the affine surface
$Y:z^{2}+f(x,y)=0$ (2)
in $\mathbb{A}^{3}$ with coordinate $(x,y,z)$. The polynomial $f$ defines a plane curve $Z(f)\subset\mathbb{A}^{2}_{x,y}$ with simple singular point $(0,0)$ of multiplicity two or three. Recall that the multiplicity of the point $(0,0)\in Z(f)$ is the lowest degree of the monomials of $f$, denoted by $\mu_{(0,0)}(f)$. Notice that in this case, $(0,0,0)$ is the only singularity of $Y$ and $(0,0)$ is the only singularity of $Z(f)$. The expression of $f$ for type $A_{n},D_{n},E_{6},E_{7},E_{8}$ can be:
$x^{2}+y^{n+1},y(x^{2}+y^{n-2}),x^{3}+y^{4},x(x^{2}+y^{3}),x^{3}+y^{5}$
respectively.
Consider the blowup $p\colon\mathrm{Bl}_{o}(\mathbb{A}^{3})\to\mathbb{A}^{3}$. Let $\iota$ be the involution of $\mathbb{A}^{3}$ sending $(x,y,z)$ to $(x,y,-z)$, which naturally extends to $\mathrm{Bl}_{o}(\mathbb{A}^{3})$. For simplicity, all restrictions and extensions of $\iota$ are still denoted by $\iota$. Clearly $\iota$ preserves the strict transform $Y_{1}\subset\mathrm{Bl}_{o}(\mathbb{A}^{3})$ of $Y$ and the exceptional plane. Hence $\iota$ also preserves the union of exceptional curve(s) in $Y_{1}$. The involution $\iota\colon Y\to Y$ is a deck transformation of the double cover $Y\to\mathbb{A}^{2}_{x,y}$. The (possibly singular) surface $Y_{1}$ also admits a double covering map to $Y_{1}/\iota$.
The following proposition characterizes the number of exceptional curve(s) arising from $p\colon Y_{1}\to Y$ and involution on the exceptional curve(s).
Proposition  2.1.
The number of exceptional curve(s) is one except for $A_{n}(n\geq 2)$. In the case of $A_{n}(n\geq 2)$, the two exceptional curves are exchanged by $\iota$. Otherwise, if $\mu_{(0,0)}f=3$ ($D_{n}(n\geq 4)$, $E_{6},E_{7},E_{8}$), then the exceptional curve is pointwisely fixed by $\iota$; if $\mu_{(0,0)}f=2$ ($A_{1}$), then $\iota$ acts on the exceptional curve as an involution.
Proposition  2.2.
The blowup $p\colon Y_{1}\to Y$ is compatible with the two involutions on $Y_{1}$ and $Y$, hence descends to a map $Y_{1}/\iota\to\mathbb{A}^{2}$ which is the blowup at $(0,0)\in\mathbb{A}^{2}$.
The above two propositions are direct from computation with affine coordinates. They imply the following:
Corollary  2.3.
If $\mu_{(0,0)}f=3$, the exceptional curve of $Y_{1}\to Y$ is pointwisely fixed by $\iota$.
The following lemma is known to experts, and we present it here for readers’ convenience. The variety $Y$ and $Y_{1}$ are normal, and we denote by $K_{Y},K_{Y_{1}}$ their canonical classes.
Lemma  2.4.
The blowup $p\colon Y_{1}\to Y$ induces $K_{Y_{1}}=p^{*}K_{Y}$.
Proof.
Denote by $A\coloneqq\mathrm{Bl}_{o}(\mathbb{A}^{3})$ and by $E$ the exceptional plane. By the adjunction formula (see [KM98, Proposition 5.73]), we have
$K_{Y_{1}}$ $=(K_{A}+Y_{1})|_{Y_{1}}$
$=(p^{*}K_{\mathbb{A}^{3}}+2E+p^{*}Y-2E)|_{Y_{1}}=p^{*}(K_{\mathbb{A}^{3}}+Y)|_{Y_{1}}=p^{*}K_{Y}.$
Thus the lemma follows. ∎
Consider a smooth projective surface $\widehat{S}$ and a reduced curve $Z\subset\widehat{S}$ with only simple curve singularities. We assume that there exists a double cover $\widehat{X}$ of $\widehat{S}$ branched along $Z$. Then $\widehat{X}$ has only ADE singularities. Denote the double cover involution of $\widehat{X}$ by $\iota$. Let $p\colon X\to\widehat{X}$ be the minimal resolution of $\widehat{X}$. The involution $\iota$ naturally extends to $X$. Denote by $S\coloneqq X/\iota$.
Proposition  2.5.
The resolution $p$ is a crepant resolution and is compatible with the involutions on $X$ and $\widehat{X}$. Hence $p$ descends to $S\to\widehat{S}$.
Proof.
The resolution $p\colon X\to\widehat{X}$ is a composition of finitely many blowups of ADE singularities. It follows from Lemma 2.4 that $p$ is a crepant resolution. By Proposition 2.2, $p$ is compatible with the two involutions. ∎
Assume that there is a singularity $q\in Z$ of ADE type $R$. The preimage of $q$ in $X$ is a union of exceptional curves, which we call the resolution graph of $q$. The following is a corollary of Proposition 2.1.
Proposition  2.6.
The Dynkin diagram of type $R$ is the dual graph of the resolution graph of $q$. The involution $\iota$ naturally acts on the Dynkin diagram of $R$. If $R=A_{1},D_{2n},E_{7},E_{8}$, the $\iota$-action on the Dynkin diagram is identity. In other cases, the $\iota$-action on the Dynkin diagram is the unique nontrivial involution of the corresponding Dynkin diagram.
Remark  2.7.
We denote by $L(R)$ the associated root lattice of type $R$. The resolution of $q$ induces a natural embedding $L(R)\hookrightarrow H^{2}(X,\mathbb{Z})$. Since the exceptional curves form a base of $L(R)$, the induced $\iota$-action on $L(R)$ is given by Propsition 2.6.
We also recall an alternative construction of the resolution. This can also be found in [BHPVdV04, III, §7]. Denote by $p\colon S_{1}\to\widehat{S}$ the blowup of $\widehat{S}$ at a singularity $q$ and by $E$ the exceptional curve. Let $\widetilde{Z}$ be the strict transform of $Z$.
Lemma  2.8.
If $\mu_{q}(Z)=2$, then $2K_{S_{1}}+\widetilde{Z}=p^{*}(2K_{\widehat{S}}+Z)$. If $\mu_{q}(Z)=3$, then $2K_{S_{1}}+(\widetilde{Z}+E)=p^{*}(2K_{\widehat{S}}+Z)$.
Proof.
We have $p^{*}Z=\widetilde{Z}+\mu_{q}(Z)E$ and $K_{S_{1}}=p^{*}K_{\widehat{S}}+E$. Then it is direct to check $2K_{S_{1}}+(\widetilde{Z}+(\mu_{q}(Z)-2)E)=2p^{*}K_{\widehat{S}}+p^{*}Z$. ∎
Lemma  2.9.
Assume that $Z$ is $2$-divisible, namely, there exists a line bundle $\mathcal{L}$ such that $\mathcal{L}^{\otimes 2}=\mathcal{O}_{\widehat{S}}(Z)$. If $\mu_{q}(Z)=2$ (respectively, $3$), then $\widetilde{Z}$ (respectively, $\widetilde{Z}+E$) is $2$-divisible.
Proof.
It is direct to check $\mathcal{O}_{S_{1}}(\widetilde{Z}+(\mu_{q}(Z)-2)E)=(p^{*}\mathcal{L}\otimes p^{*}\omega_{\widehat{S}}\otimes\omega_{S_{1}}^{-1})^{\otimes 2}$. ∎
Under the assumption of Lemma 2.9, there exist double covers $\widehat{\pi}\colon\widehat{X}\to\widehat{S}$ (respectively $\pi_{1}\colon X_{1}\to S_{1}$) branched along $Z$ (respectively $\widetilde{Z}+(\mu_{q}(Z)-2)E$). We denote by $Z_{1}=\widetilde{Z}+(\mu_{q}(Z)-2)E$.
Proposition  2.10.
There exists a natural birational morphism $\widetilde{p}\colon X_{1}\to\widehat{X}$ such that $\widehat{\pi}\circ\widetilde{p}=p\circ\pi_{1}$ and $\widetilde{p}^{*}K_{\widehat{X}}\cong K_{X_{1}}$.
Proof.
Since $p\colon S_{1}-p^{-1}(q)\to\widehat{S}-q$ is an isomorphism, it naturally lifts to an isomorphism $\widetilde{p}\colon X_{1}-\pi_{1}^{-1}(E)\cong\widehat{X}-\pi^{-1}(q)$ satisfying $\widehat{\pi}\circ\widetilde{p}=p\circ\pi_{1}$. Moreover, $\widetilde{p}$ extends to a morphism $X_{1}\to\widehat{X}$. Notice $2K_{X_{1}}=\pi_{1}^{*}(2K_{S_{1}}+Z_{1})$ and $2K_{\widehat{X}}=\widehat{\pi}^{*}(2K_{\widehat{S}}+Z)$. By Lemma 2.8, $\widetilde{p}^{*}K_{\widehat{X}}\cong K_{X_{1}}$. ∎
The blowup gives a new pair $(S_{1},Z_{1})$ from $(\widehat{S},Z)$. We iterate this step until we get a smooth $Z_{n}$ (the iteration must end in finite steps). Denote by $(\overline{S},\overline{Z})\coloneqq(S_{n},Z_{n})$. The smooth curve $\overline{Z}$ is a disjoint union of irreducible smooth curves.
Proposition  2.11.
The rational surface $\overline{S}$ admits a smooth double cover $\overline{X}$ branched along $\overline{Z}$. We have $\overline{X}\cong X$.
Proof.
By Proposition 2.10, $\overline{X}$ admits a birational morphism $\overline{p}\colon\overline{X}\to\widehat{X}$ such that $\overline{p}^{*}K_{\widehat{X}}\cong K_{\overline{X}}$. Since $X$ is minimal, $\overline{p}$ factors through $\rho\colon\overline{X}\to X$, and $\rho^{*}K_{X}\cong K_{\overline{X}}$. This implies $\overline{X}\cong X$. ∎
2.3. K3 Surfaces Associated with ADE Sextic Curves
Let $F$ be a homogeneous sextic polynomial of three variables, and $Z(F)$ be the associated sextic curve in $\mathbb{P}^{2}$. We assume that $Z(F)$ has only simple singularities. Let $\widehat{X}$ be the double cover of $\mathbb{P}^{2}$ branched along $Z(F)$. Then $K_{\widehat{X}}$ is trivial and we call $\widehat{X}$ an ADE K3 surface.
By the construction in §2.2, we have a pair $(S_{F},Z_{F})$ from $(\mathbb{P}^{2},Z(F))$. Denote by $X_{F}$ the double cover of $S_{F}$ branched over $Z_{F}$. From Proposition 2.11 we know:
Proposition  2.12.
The surface $X_{F}$ is the minimal resolution of $\widehat{X}$. Moreover, $X_{F}$ is a smooth K3 surface.
Fix certain $F\in\mathcal{V}_{T}$ for a given singular type $T$. Denote by $p=p_{F}\colon X_{F}\to S_{F}$ the branched double covering. Let $\iota=\iota_{F}$ be the involution on $X_{F}$ induced by the double covering.
Denote by $\Lambda_{F}$ the lattice $H^{2}(X_{F},\mathbb{Z})$. Let $H_{F}\in H^{2}(X_{F},\mathbb{Z})$ be the pullback of the hyperplane class in $\mathbb{P}^{2}$. Let $l_{T}(R)$ be the number of singularities of type $R$ that appear in the singular K3 surface associated with an element in $\mathcal{V}_{T}$. Denote by $L_{F}=\oplus_{R}L(R)^{l_{T}(R)}$, where $R$ runs through all irreducible root systems of ADE type. Let $P_{F}$ be the primitive hull of $L_{F}\oplus\langle H_{F}\rangle$ in $H^{2}(X_{F},\mathbb{Z})$. Notice that $P_{F}$ is a primitive sublattice of the Picard lattice $\mathrm{Pic}(X_{F})$. The induced action of $\iota$ on $\Lambda_{F}$ is still denoted by $\iota$. This action naturally preserves $\mathrm{Pic}(X_{F})$, $L_{F}$ and $P_{F}$. By Proposition 2.5, Proposition 2.6 and Remark 2.7, we conclude the following:
Proposition  2.13.
The pullback $p^{*}\colon H^{2}(S,\mathbb{Z})\to H^{2}(X,\mathbb{Z})$ induces a lattice saturation $H^{2}(S,\mathbb{Z})\hookrightarrow\langle H_{F}\rangle\oplus L_{F}^{\iota}$. Here $L_{F}^{\iota}$ is the $\iota$-invariant sublattice.
Corollary  2.14.
We have $\Lambda_{F}^{\iota}=P_{F}^{\iota}$. In particular, $\iota$ acts as $-1$ on $P_{F}^{\perp}$, where $P_{F}^{\perp}$ denotes the orthogonal complement of $P_{F}$ in $\Lambda_{F}$.
Proof.
By Lemma 5.1 (a general topological result, whose proof does not depend on the rest of the paper), $p^{*}$ induces a saturation $H^{2}(S,\mathbb{Z})\hookrightarrow\Lambda_{F}^{\iota}$. Since $\Lambda_{F}^{\iota}$ is primitive in $\Lambda_{F}$, there exists a saturation $\langle H_{F}\rangle\oplus L_{F}^{\iota}\hookrightarrow\Lambda_{F}^{\iota}$. Since $P_{F}$ is the primitive hull of $\langle H_{F}\rangle\oplus L_{F}$ in $\Lambda_{F}$, we have $\Lambda_{F}^{\iota}\subset P_{F}$. Thus $\Lambda_{F}^{\iota}=P_{F}^{\iota}$. ∎
Remark  2.15.
For a nontrivial singular type $T$, the associated K3 surfaces $X_{F}$ always admit natural elliptic fibration structures. Explicitly, consider the pencil of lines on $\mathbb{P}^{2}$ through any singularity of $Z(F)$. The preimage of a generic such line in $X_{F}$ is an elliptic curve. Therefore, the K3 surface $X_{F}$ admits an elliptic fibration. Note that this elliptic fibration may not admit any section (but always admit a multi-section).
3. Occult Period Map and Global Torelli
In this section, we study the Hodge structures of the associated K3 surfaces $X_{F}$ for $F\in\mathcal{V}_{T}$, and construct the corresponding period map.
3.1. Period Domain, Monodromy Group and Occult Period Map
We first define the period domain and monodromy group for a fixed singular type $T$.
Denote by $Q_{F}=P_{F}^{\perp}$ the orthogonal complement of $P_{F}$ in $\Lambda_{F}$. Then $P_{F}$ and $Q_{F}$ are sublattices of $H^{2}(X_{F},\mathbb{Z})$ with signatures $(1,\sum_{R}l_{T}(R)\mathrm{rank}(R))$ and $(2,19-\sum_{R}l_{T}(R)\mathrm{rank}(R))$ respectively.
Definition  3.1 (Period Domain).
The space $\mathbb{P}\{x\in(Q_{F})_{\mathbb{C}}\big{|}x\cdot x=0,x\cdot\overline{x}>0\}$ has two connected components that are interchanged by complex conjugation. Denote by $\mathbb{D}(Q_{F})$ the connected component that contains $H^{2,0}(X_{F})$. This is the period domain for surfaces $X_{F}$.
Let $\Delta_{F}$ be the set of exceptional curves arising from the resolution of singularities, which is a base of $L_{F}$, see §2.3. Denote by $\mathrm{O}(\Lambda_{F},\Delta_{F},H_{F})$ the group of automorphisms $f\in\mathrm{O}(\Lambda_{F})$ such that $f(\Delta_{F})=\Delta_{F}$ and $f(H_{F})=H_{F}$. An automorphism of $\Lambda_{F}$ which preserves $\Delta_{F}$ and $H_{F}$ necessarily preserves $P_{F}$ and, consequently, induces an automorphism of $Q_{F}$. Define $\widetilde{\Gamma}_{F}$ to be the image of the homomorphism
$\mathrm{O}(\Lambda_{F},\Delta_{F},H_{F})\to\mathrm{O}(Q_{F}).$
Let $\Gamma_{F}$ be the subgroup of $\widetilde{\Gamma}_{F}$ that preserves $\mathbb{D}(Q_{F})$. Then $[\widetilde{\Gamma}_{F}:\Gamma_{F}]\leq 2$, with equality holds if and only if there exists an element in $\widetilde{\Gamma}_{F}$ interchanging the two components. By Corollary 2.14, the $\iota_{F}$ acts as $-1$ on $Q_{F}$. Thus $-1\in\Gamma_{F}$ holds for all singular types. The following proposition is direct from lattice theory.
Proposition  3.2.
Any element $g\in\mathrm{O}(Q_{F})$ that acts trivially on the discriminant group $A_{Q_{F}}$ belongs to $\widetilde{\Gamma}_{F}$. In particular, $\Gamma_{F}$ is of finite index in $\mathrm{O}(Q_{F})$, hence an arithmetic group acting on $\mathbb{D}(Q_{F})$.
From now on, we choose a base point $F_{0}\in\mathcal{V}_{T}$, and denote by $\Gamma_{T}=\Gamma_{F_{0}}$, $\widetilde{\Gamma}_{T}=\widetilde{\Gamma}_{F_{0}}$ and $\mathbb{D}_{T}=\mathbb{D}(Q_{F_{0}})$.
For $F\in\mathcal{V}_{T}$, choose a path $\gamma$ connecting $F$ and $F_{0}$. The path $\gamma$ induces a diffeomorphism from $X_{F}$ to $X_{F_{0}}$ together with an isomorphism
$\gamma^{*}\colon(\Lambda_{F},\Delta_{F},H_{F},P_{F},Q_{F})\cong(\Lambda_{F_{0}},\Delta_{F_{0}},H_{F_{0}},P_{F_{0}},Q_{F_{0}}).$
The line $\gamma^{*}(H^{2,0}(X_{F}))$ represents a point in $\mathbb{D}_{T}=\mathbb{D}(Q_{F_{0}})$.
Next we define the period map for $X_{F}$. Choose any two paths $\gamma$, $\gamma^{\prime}$ in $\mathcal{V}_{T}$ connecting $F$ and $F_{0}$. Then $\gamma^{\prime*}\circ(\gamma^{*})^{-1}$ is an automorphism of $(\Lambda_{F_{0}},\Delta_{F_{0}},H_{F_{0}},P_{F_{0}},Q_{F_{0}})$, which induces an element in $\Gamma_{T}$. Therefore, we have an analytic map
$\mathscr{P}_{T}\colon\mathcal{V}_{T}\to\Gamma_{T}\backslash\mathbb{D}_{T},$
which descends to
$\mathscr{P}_{T}\colon\mathcal{M}_{T}\to\Gamma_{T}\backslash\mathbb{D}_{T}.$
Following [KR12], we call $\mathscr{P}_{T}$ the occult period map for sextic curves of type $T$.
3.2. Global Torelli
We continue with the notation in §3.1. Based on the global Torelli theorem for K3 surfaces [BR75], we prove the following global Torelli theorem for occult period maps of singular sextic curves.
Theorem  3.3.
For any type $T$, the period map $\mathscr{P}_{T}\colon\mathcal{M}_{T}\to\Gamma_{T}\backslash\mathbb{D}_{T}$ is injective.
Proof.
Taking $F$, $F^{\prime}\in\mathcal{V}_{T}$ such that $\mathscr{P}_{T}(F)=\mathscr{P}_{T}(F^{\prime})$, we aim to show that $Z(F)$ is linearly isomorphic to $Z(F^{\prime})$. We first show that the K3 surfaces $X_{F}$ and $X_{F^{\prime}}$ are isomorphic. Choose two paths $\gamma_{F}$ and $\gamma_{F^{\prime}}$ connecting $F$ and $F^{\prime}$ to $F_{0}$. There exists an element $g\in\Gamma_{T}$ such that $g\cdot\gamma_{F}^{*}(H^{2,0}(X_{F}))=\gamma_{F^{\prime}}^{*}(H^{2,0}(X_{F^{\prime}}))$. By the definition of $\Gamma_{T}$, there exists $\widetilde{g}\in\mathrm{O}(\Lambda_{F_{0}},\Delta_{F_{0}},H_{F_{0}})$ whose restriction to $Q_{F_{0}}$ is $g$. Then the isomorphism
$\phi=(\gamma_{F^{\prime}}^{*})^{-1}\circ\widetilde{g}\circ\gamma_{F}^{*}\colon(\Lambda_{F},\Delta_{F},H_{F})\cong(\Lambda_{F^{\prime}},\Delta_{F^{\prime}},H_{F^{\prime}})$
maps $H^{2,0}(X_{F})$ to $H^{2,0}(X_{F^{\prime}})$.
Choose an element $\beta_{F}$ in the Weyl chamber of $L_{F}$ relative to $\Delta_{F}$ and let $\beta_{F^{\prime}}=\phi(\beta_{F})$. For a smooth rational curve $D$ in $X_{F}$, we have either $[D]\cdot H_{F}$ a positive integer or $[D]\in\Delta_{F}$. The latter case implies $[D]\cdot\beta_{F}>0$. Thus there exists a positive integer $c$ such that $\alpha_{F}=\beta_{F}+cH_{F}$ has a positive intersection with any such $[D]$. By Nakai–Moishezon criterion for K3 surfaces, $\alpha_{F}$ is ample. If $c$ is large enough, we may ask $\alpha_{F^{\prime}}=\beta_{F^{\prime}}+cH_{F^{\prime}}$ also ample. Then the isomorphism $\phi\colon\Lambda_{F}\to\Lambda_{F^{\prime}}$ is a Hodge isometry that identifies the two ample classes $\alpha_{F}$ and $\alpha_{F^{\prime}}$. By the global Torelli theorem for K3 surfaces [BR75], there exists a unique isomorphism $\Phi\colon X_{F}\to X_{F^{\prime}}$ induces $\phi$.
We finally aim to construct an isomorphism of $Z(F)$ to $Z(F^{\prime})$. Notice that $\phi$ is compatible with the deck transformations $\iota_{F}$ and $\iota_{F^{\prime}}$. By the faithfulness of the induced action of automorphisms of K3 surfaces on the middle cohomology [LP81], $\Phi$ is compatible with $\iota_{F}$ and $\iota_{F^{\prime}}$. Hence $\Phi$ induces isomorphisms $\mathrm{Fix}(\iota_{F})\cong\mathrm{Fix}(\iota_{F^{\prime}})$ and $S_{F}\cong S_{F^{\prime}}$. Furthermore, $\Phi$ induces isomorphism between linear systems of $H_{F}$ and $H_{F^{\prime}}$. Therefore, $\Phi$ descends to a linear isomorphism $(\mathbb{P}^{2},Z(F))\cong(\mathbb{P}^{2},Z(F^{\prime}))$. The injectivity of $\mathscr{P}_{T}\colon\mathcal{M}_{T}\to\Gamma_{T}\backslash\mathbb{D}_{T}$ follows. ∎
3.3. Image of the Period Map $\mathscr{P}_{T}$
Next we characterize the image of the period map $\mathscr{P}_{T}$. Many of the results in this section are essentially a reinterpretation of [Ura88, Theorem 1.16] and [Deg08, Theorem 3.4.1]. Our description focuses on the level of moduli spaces and arithmetic quotients.
Since we fix the base point $F_{0}$, we omit the subscript $F_{0}$ for $\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}},L_{F_{0}},P_{F_{0}},Q_{F_{0}}$ throughout this section. Let $\mathcal{H}_{T}$ be the union of $r^{\perp}\subset\mathbb{D}_{T}$ with $r$ running through all roots $r$ such that $r\perp H$ and $r\notin L$.
Theorem  3.4.
For any type $T$, the period map $\mathscr{P}_{T}\colon\mathcal{M}_{T}\to\Gamma_{T}\backslash\mathbb{D}_{T}$ induces an isomorphism $\mathcal{M}_{T}\cong\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$.
The orthogonal complement $H^{\perp}$ of $H$ in $\Lambda$ has signature $(2,19)$ and defines a $19$-dimensional type IV domain $\mathbb{D}(H^{\perp})$. There is a natural inclusion $\mathbb{D}_{T}\subset\mathbb{D}(H^{\perp})$.
Lemma  3.5.
For any $\omega\in\mathbb{D}_{T}-\mathcal{H}_{T}$, there does not exist $u\in\omega^{\perp}_{\Lambda}$ such that $u^{2}=0$ and $u\cdot H=1$.
Proof.
Suppose there exists an element $u$ satisfying the requirement. We observe that $(2u-H)^{2}=-2$ and $(2u-H)\cdot H=0$, thus $2u-H$ is a root in $\langle H,\omega\rangle^{\perp}_{\Lambda}$. By the choice of $\omega$, we have $2u-H\in L$, then $u\in P\subset\mathrm{Pic}(X_{F_{0}})$. Hence $\mathrm{Pic}(X_{F_{0}})$ contains $u$ such that $u^{2}=0$ and $u\cdot H=1$. By [Ura88, Proposition 1.7], this contradicts the fact that a line bundle on $X_{F_{0}}$ representing $H$ defines a surjective morphism $X_{F_{0}}\to\mathbb{P}^{2}$ of degree two. ∎
Denote by $\omega_{0}\coloneqq H^{2,0}(X_{F_{0}})$. Take any $\omega\in\mathbb{D}_{T}-\mathcal{H}_{T}$. By surjectivity of the period map for marked complex K3 surfaces, there exists a K3 surface $X_{\omega}$ with a marking $\phi_{\omega}\colon H^{2}(X_{\omega},\mathbb{Z})\to\Lambda$ such that $\phi_{\omega}(H^{2,0}(X_{\omega}))=\omega$. From now on, we fix $\omega$ throughout the proof. The following lemma is essentially proved by Urabe [Ura88, Theorem 1.16]:
Lemma  3.6.
The K3 surface $X_{\omega}$ is the minimal resolution of the double cover of $\mathbb{P}^{2}$ branched along a singular sextic curve $Z_{\omega}$. The singular sets of $Z_{\omega}$ is determined by $\Delta$.
Proof.
By [Huy16, Chapter 8, §2, Corollary 2.9], there exist rational curves $C_{1},\cdots,C_{n}$ on $X_{\omega}$ such that
$s_{[C_{1}]}\circ\cdots\circ s_{[C_{n}]}(\phi_{\omega}^{-1}H)$
is a nef class with self-intersection number $2$, where $[C_{i}]$ is a root in $\mathrm{Pic}(X_{\omega})$ and $s_{[C_{i}]}$ denotes the refleciton with respect to the root $[C_{i}]$. Since $s_{[C_{i}]}$ acts trivially on $A_{\mathrm{Pic}(X_{\omega})}$, it naturally extends to an isometry $\widetilde{s_{[C_{i}]}}$ of $H^{2}(X_{\omega},\mathbb{Z})$ that acts trivally on $\mathrm{Pic}(X_{\omega})^{\perp}$. Denote by
$\phi^{\prime}_{\omega}\coloneqq\phi_{\omega}\circ\widetilde{s_{[C_{n}]}}\circ\cdots\circ\widetilde{s_{[C_{1}]}}.$
The new marking $\phi^{\prime}_{\omega}$ of $X_{\omega}$ satisfies that $\phi^{\prime-1}_{\omega}H$ is nef and $\phi^{\prime}_{\omega}(H^{2,0}(X_{\omega}))=\omega$. By [Ura88, Proposition 1.7] and Lemma 3.5, the complete linear system of $\phi^{\prime-1}_{\omega}H$ defines a surjective morphism $\pi_{\omega}\colon X_{\omega}\to\mathbb{P}^{2}$ of degree $2$.
By the choice of $\omega$, the root lattice of $\{x\in H^{\perp}\big{|}x\perp\omega\}$ equals to $L$. (Note that $\omega$ not only determines the root lattice $L$ up to isomorphism, but also determines an embedding of $L$ into $H^{\perp}\subset\Lambda$.) We first show that all rational curves contracted by $\pi_{\omega}$ form a base of $\phi^{\prime-1}_{\omega}L$. Let $C$ be a connected curve in $X$ that is contracted by $\pi_{\omega}$. This is equivalent to say that $C\cdot\phi^{\prime-1}_{\omega}H=0$. By Hodge index theorem, $C^{2}<0$, hence $C^{2}=-2$. This implies $[C]\in\phi^{\prime-1}_{\omega}L$. Moreover, $C$ is a connected $\mathbb{P}^{1}$-graph whose dual graph is a Dynkin diagram of certain ADE type. Conversely, any rational curve in $X$ that defines a root in $L$ is contracted by $\pi_{\omega}$, since it is orthogonal to $\phi^{\prime-1}_{\omega}H$. The cohomology classes of all these rational curves generate the root lattice $\phi^{\prime-1}_{\omega}L$. Hence they form a base of $\phi^{\prime-1}_{\omega}L$.
This concludes that $\pi_{\omega}$ factor through a birational morphism $X_{\omega}\to\widehat{X}_{\omega}$ that contracts all rational curves orthogonal to $\phi^{\prime-1}_{\omega}H$, where these rational curves form the unique effective base of $\phi^{\prime-1}_{\omega}L$. Thus the base is exactly $\phi^{\prime-1}_{\omega}\Delta$. The normal surface $\widehat{X}_{\omega}$ is a singular K3 surface with ADE singularities, where the singularities are the image of rational curves in $\phi^{\prime-1}_{\omega}\Delta$. Since $\pi_{\omega}$ is surjective, the induced map $p_{\omega}\colon\widehat{X}_{\omega}\to\mathbb{P}^{2}$ is also surjective. Hence $p_{\omega}$ is a branched double covering. By [BHPVdV04, III, §7, Theorem 7.2], the branched locus of $p_{\omega}$ is a sextic curve $Z_{\omega}$ on $\mathbb{P}^{2}$. Therefore, the map $\pi_{\omega}$ is a double cover branched along a sextic curve composed of the contraction of $(-2)$-curves on $X_{\omega}$, and the singularities of the sextic curve $Z_{\omega}$ are the images of $(-2)$-curves in $\phi^{\prime-1}_{\omega}\Delta$. ∎
The following lemma is essentially proved by Degtyarev [Deg08, Theorem 3.4.1]. We provide an alternative proof for the reader’s convenience.
Lemma  3.7.
The singular sextic curve $Z_{\omega}$ is equisingular deformation equivalent to $Z(F_{0})$, namely, $Z_{\omega}\in\mathbb{P}\mathcal{V}_{T}$.
Proof.
We choose arbitrarily a path $\gamma_{\omega}$ in $\mathbb{D}_{T}-\mathcal{H}_{T}$ connecting $\omega$ and $\omega_{0}$. For any $\omega^{\prime}\in\gamma_{\omega}$, by the proof of Lemma 3.6, there exists a marking $\phi_{\omega^{\prime}}$ of $X_{\omega^{\prime}}$ such that $\phi_{\omega^{\prime}}^{-1}H$ is nef. By [Ura88, Proposition 1.7] and [Huy16, Chapter 2, §3, Corollary 3.14], any line bundle $\mathcal{O}(C_{\omega^{\prime}})$ representing the cohomology class $\phi_{\omega^{\prime}}^{-1}H$ is base point free, where $C_{\omega^{\prime}}$ is a nef divisor in $X_{\omega^{\prime}}$. By Bertini’s theorem, a generic member in the complete linear system $|C_{\omega^{\prime}}|$ is a smooth genus two curve. We may assume that $C_{\omega^{\prime}}$ is a smooth genus two curve. Denote by $\mathcal{E}_{\omega^{\prime}}\coloneqq\{E_{r}\,|\,r\in\phi_{\omega^{\prime}}^{-1}\Delta\}$ be the set of exceptional curves in $X_{\omega^{\prime}}$ representing $\phi_{\omega^{\prime}}^{-1}\Delta$.
The universal deformation of $X_{\omega^{\prime}}$ in the moduli space of complex K3 surfaces induces a family of K3 surfaces over a contractible complex neighborhood $D$ of $\omega^{\prime}$ in $\mathbb{D}_{T}-\mathcal{H}_{T}$, by the local Torelli theorem. The marking $\phi_{\omega^{\prime}}$ of $X_{\omega^{\prime}}$ canonically induces markings for all fibers $H^{2}(X_{\eta},\mathbb{Z})\cong H^{2}(X_{\omega^{\prime}},\mathbb{Z})\cong\Lambda$, $\eta\in D$. Since $\eta^{\perp}_{\Lambda}$ contains $H$ and $\Delta$, $\phi_{\omega^{\prime}}^{-1}H$ and $\phi_{\omega^{\prime}}^{-1}\Delta$ lifts to algebraic classes over $D$. Due to [Nis24, Page 363], any curve on a K3 surface is semiregular. By [Blo72, Theorem 7.1] and [Art68, Theorem 1.5(i)], there exists a complex neighborhood $D_{\omega^{\prime}}\subset D$ of $\omega^{\prime}$ in $\mathbb{D}_{T}-\mathcal{H}_{T}$ such that for any $\eta\in D_{\omega^{\prime}}$, $C_{\omega^{\prime}}$ and $\mathcal{E}_{\omega^{\prime}}$ deform to smooth algebraic curves $C_{\eta}$ and $\mathcal{E}_{\eta}$ on $X_{\eta}$, where $[C_{\eta}]=\phi_{\eta}^{-1}H$.
Denote by $\mathcal{X}$ the family of K3 surfaces over $D_{\omega^{\prime}}$. For any $\eta\in D_{\omega^{\prime}}$, the complete linear system of $\mathcal{O}(C_{\eta})$ defines a morphism $X_{\eta}\to\mathbb{P}H^{0}(\mathcal{O}(C_{\eta}))^{\vee}\cong\mathbb{P}^{2}$ of degree two. There exists a line bundle $\mathcal{L}$ on $\mathcal{X}$ such that $\mathcal{L}_{\eta}=\mathcal{O}(C_{\eta})$ for any $\eta\in D_{\omega^{\prime}}$. Denote by $\lambda\colon\mathcal{L}\to\mathcal{X}$ the natural projection. We have a relative projective morphism $\mathcal{X}\to\mathbb{P}R^{0}\lambda_{*}\mathcal{L}^{\vee}$ over $D_{\omega^{\prime}}$ defined by the relative linear system $|\mathcal{L}|$. Note that $\mathbb{P}R^{0}\lambda_{*}\mathcal{L}^{\vee}$ is a trivial family of $\mathbb{P}^{2}$ over $D_{\omega^{\prime}}$. We can go through the process of Lemma 3.6 relatively over $D_{\omega^{\prime}}$, then obtain a family of singular sextic curves $\mathcal{Z}\subset\mathbb{P}R^{0}\lambda_{*}\mathcal{L}^{\vee}$ sharing the same type of singularities. Then any two sextic curves in this family are equisingular deformaiton equivalent.
The path $\gamma_{\omega}$ is covered by finite complex neighborhoods in $\mathbb{D}_{T}-\mathcal{H}_{T}$, and this concludes that $(Z_{\omega},\mathbb{P}^{2})$ is equisingular deformation equivalent to $(Z(F_{0}),\mathbb{P}^{2})$. ∎
Then the sextic curve $Z_{\omega}$ can be written as $Z(F)$ for some sextic polynomial $F\in\mathcal{V}_{T}$. The associated K3 surface $X_{\omega}$ is the same as $X_{F}$, which we construct in §2.3.
Lemma  3.8.
The occult period map $\mathscr{P}_{T}$ maps $Z(F)$ to $[\omega]$ in $\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$.
Proof.
We choose arbitrarily a path $\gamma_{F}$ connecting $Z(F)$ and $Z(F_{0})$ in $\mathbb{P}\mathcal{V}_{T}$. The path $\gamma_{F}$ canonically induces a marking $\gamma_{F}^{*}\colon H^{2}(X_{F},\mathbb{Z})\to\Lambda$. The induced cohomology class $H_{F}\coloneqq(\gamma_{F}^{*})^{-1}H\in H^{2}(X_{F},\mathbb{Z})$ is independent of the choice of the path $\gamma_{F}$, which is equal to the pullback of the hyperplane class in $\mathbb{P}^{2}$. By the proof of Lemma 3.6, there exists a marking $\phi_{\omega}$ of $X_{\omega}=X_{F}$ such that $\phi_{\omega}^{-1}H=H_{F}$ and $\phi_{\omega}^{-1}\Delta=\Delta_{F}$. Consider $g\coloneqq\gamma_{F}^{*}\circ\phi_{\omega}^{-1}\in\mathrm{O}(\Lambda)$, it preserves $H$ and $\Delta$. Thus the restriction $g|_{Q}$ belongs to $\widetilde{\Gamma}_{T}$. It is direct $g\cdot\omega=\gamma_{F}^{*}H^{2,0}(X_{F})$. Both $\omega$ and $\gamma_{F}^{*}H^{2,0}(X_{F})$ lies in $\mathbb{D}_{T}$, hence $g$ sends $\mathbb{D}_{T}$ to $\mathbb{D}_{T}$. It follows that $g|_{Q}\in\Gamma_{T}$. We conclude that $\mathscr{P}_{T}$ maps $Z(F)$ to $[g\cdot\omega]=[\omega]$ in $\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$. ∎
Proof of Theorem 3.4.
It follows directly from Lemma 3.6, Lemma 3.7 and Lemma 3.8. ∎
Corollary  3.9.
For any generic $F\in\mathcal{V}_{T}$ we have $\mathrm{Pic}(X_{F})=P_{F}$.
Proof.
By Theorem 3.4, the period map $\mathscr{P}_{T}\colon\mathcal{M}_{T}\to\Gamma_{T}\backslash\mathbb{D}_{T}$ is open. For a generic $F\in\mathcal{V}_{T}$, the period of $F$ is a general point in $\mathbb{D}_{T}$. Thus the Picard lattice of $X_{F}$ is $P_{F}$. ∎
Remark  3.10.
The hyperplane arrangement $\mathcal{H}_{T}$ usually has infinite irreducible components, but the quotient $\Gamma_{T}\backslash\mathcal{H}_{T}$ is, in fact, a divisor of $\Gamma_{T}\backslash\mathbb{D}_{T}$. This can be concluded from [BB66], [Bor72] and Theorem 3.4.
Remark  3.11.
Many moduli spaces of sextic curves with specified singularities and group actions are closely related to Deligne–Mostow theory [DM86]. For example, Dolgachev, van Geemen and Kondō [DvGK05] studied moduli space of pairs consisting of a smooth cubic surface and a line on it. Such pairs are naturally related to pairs consisting of a smooth quintic curve and a line that admits a $\mu_{3}$-symmetry. The moduli space corresponds to the case $({1\over 6},{1\over 6},{1\over 3},{1\over 3},{1\over 3},{1\over 3},{1\over 3})$ in [Mos88]. In [ZZ24] the last two authors studied irreducible sextic curves with a $D_{4}$-singularity and $\mu_{3}$-symmetry, and related the moduli space to case $({1\over 6},{1\over 6},{1\over 6},{1\over 6},{1\over 6},{1\over 6},{1\over 3},{1\over 3},{1\over 3})$ in [Mos88]. The cases of nodal sextic curves with symmetries have been studied in [YZ23, §5].
4. GIT Compactification and Arithmetic Compactification
4.1. K3 Surfaces of Degree $2$
We first review the classical results of moduli space of sextic curves via Hodge theory and compactifications. See [Sha80] and Looijenga.
Let $V$ be a $3$-dimensional complex vector space. A (plane) sextic curve is the zero locus of an element of $\mathrm{Sym}^{6}V^{\vee}$ in $\mathbb{P}V$. By definition, $\mathbb{P}\mathrm{Sym}^{6}V^{\vee}$ is the space of sextic curves. Let $\mathbb{P}(\mathrm{Sym}^{6}V^{\vee})^{\mathrm{sm}}$ be the subset of $\mathbb{P}\mathrm{Sym}^{6}V^{\vee}$ consisting of smooth sextic curves. We define $\mathcal{M}$ as the GIT quotient $\mathrm{SL}(V)\backslash\!\!\backslash\mathbb{P}(\mathrm{Sym}^{6}V^{\vee})^{\mathrm{sm}}$. Denote by $\overline{\mathcal{M}}$ the GIT compactification of $\mathcal{M}$, and $\mathcal{M}^{\mathrm{ADE}}\subset\overline{\mathcal{M}}$ the subspace consisting of sextic curves with at worst simple singularities.
We briefly recall the definition of period domain and arithmetic group of the moduli space of polarized K3 surfaces of degree $2d$. Let $\Lambda\coloneqq U^{3}\oplus E_{8}^{2}$ be the K3 lattice. Let $l\coloneqq e_{1}+df_{1}$, where $e_{1},f_{1}$ is the standard basis of the first copy of $U$, and $l^{2}=2d$. Denote by $\Lambda_{d}$ for $l^{\perp}$ in $\Lambda_{K3}$. The subset of $\mathbb{P}(\Lambda_{d,\mathbb{C}})$ defined by $z\cdot z=0$ and $z\cdot\bar{z}>0$ has two connected components that are interchanged by complex conjugation. Denote by $\mathbb{D}_{d}$ and $\overline{\mathbb{D}}_{d}$ the two components. Define a subgroup of $\mathrm{O}(\Lambda)$
$\widetilde{\Gamma}_{d}\coloneqq\{g|_{\Lambda_{d}}\,\big{|}\,g\in\mathrm{O}(\Lambda),\,g(e_{1}+df_{1})=e_{1}+df_{1}\}.$
Equivalently, $\widetilde{\Gamma}_{d}$ is isomorphic to the subgroup of $\mathrm{O}(\Lambda_{d})$ which consists of elements that act trivially on $A_{\Lambda_{d}}$. Moreover, $\widetilde{\Gamma}_{d}$ is an arithmetic subgroup of $\mathrm{O}(\Lambda_{d})$. Let $\Gamma_{d}$ denote the index two subgroup of $\widetilde{\Gamma}_{d}$ that preserves $\mathbb{D}_{d}$. By Baily-Borel, the arithmetic quotient $\Gamma_{d}\backslash\mathbb{D}_{d}$ is a normal quasi-projective variety.
For a $Z\in\mathbb{P}(\mathrm{Sym}^{6}V^{\vee})^{\mathrm{sm}}$, the double cover of $\mathbb{P}V$ branched along $Z$ is a smooth K3 surface $X$. Denote by $H\in H^{2}(X,\mathbb{Z})$ the pullback of the hyperplane class of $\mathbb{P}V$, which is a degree $2$ polarization of $X$. Then the period map for polarized K3 surfaces of degree $2$ (for example, see [Huy16, Chapter 6, §4.1]) induces a period map $\mathscr{P}\colon\mathbb{P}(\mathrm{Sym}^{6}V^{\vee})^{\mathrm{sm}}\to\Gamma_{1}\backslash\mathbb{D}_{1}$. The period map $\mathscr{P}$ is $\mathrm{SL}(V)$-equivariant, hence descends to
$\mathscr{P}\colon\mathcal{M}\to\Gamma_{1}\backslash\mathbb{D}_{1}.$
We have two $\Gamma_{1}$-invariant hyperplane arrangements $\mathcal{H}_{\Delta}$ and $\mathcal{H}_{\infty}$ in $\mathbb{D}_{1}$, which correspond to two $\Gamma_{1}$-orbits of roots in $\Lambda_{1}$. The roots defining $\mathcal{H}_{\Delta}$ (respectively, $\mathcal{H}_{\infty}$) has divisibility $1$ (respectively, $2$). We have the Looijenga compactification of $\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\infty})$, denoted by $\overline{\Gamma_{1}\backslash\mathbb{D}_{1}}^{\mathcal{H}_{\infty}}$, see [Loo03]. From [Sha80] and [Loo03], we have
Theorem  4.1 (Shah, Looijenga).
The period map $\mathscr{P}$ is an algebraic open embedding whose image is $\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\Delta}-\mathcal{H}_{\infty})$. Moreover, $\mathscr{P}$ extends to an isomorphism $\mathcal{M}^{\mathrm{ADE}}\cong\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\infty})$, and further to ¯$\overline{\mathcal{M}}\cong\overline{\Gamma_{1}\backslash\mathbb{D}_{1}}^{\mathcal{H}_{\infty}}$.
4.2. Identification of GIT Compactification and Looijenga Compactification
In this section, we describe the relationship between GIT compactifications and Looijenga compactifications. We fix a singular type $T$.
4.2.1. $\mathcal{M}_{T}$ and $\overline{\mathcal{M}}$
We first describe the relation between $\mathcal{M}_{T}$ and $\overline{\mathcal{M}}$.
By definition, we have $\mathcal{M}_{T}\subset\overline{\mathcal{M}}$. Let $\mathcal{M}_{T}^{\mathrm{sst}}\coloneqq\mathrm{SL}(3)\backslash\!\!\backslash\mathbb{P}\mathcal{V}_{T}^{\mathrm{sst}}$, which is the closure of $\mathcal{M}_{T}$ in $\overline{\mathcal{M}}$. Define $\widetilde{\mathbb{P}\mathcal{V}}_{T}^{\mathrm{sst}}$ be the normalization of $\mathbb{P}\mathcal{V}_{T}^{\mathrm{sst}}$. Then the corresponding GIT quotient ¯$\overline{\mathcal{M}}_{T}\coloneqq\mathrm{SL}(3)\backslash\!\!\backslash\widetilde{\mathbb{P}\mathcal{V}}_{T}^{\mathrm{sst}}$ is the normalization of $\mathcal{M}_{T}^{\mathrm{sst}}$. In conclusion, we have
Proposition  4.2.
The natural inclusion $\mathcal{M}_{T}\subset\overline{\mathcal{M}}$ induces a generically injective finite morphism $j_{T}\colon\overline{\mathcal{M}}_{T}\to\overline{\mathcal{M}}$, which is the normalization onto its image.
4.2.2. $\Gamma_{T}\backslash\mathbb{D}_{T}$ and $\Gamma_{1}\backslash\mathbb{D}_{1}$
We next describe the relation between $\Gamma_{T}\backslash\mathbb{D}_{T}$ and $\Gamma_{1}\backslash\mathbb{D}_{1}$.
Let $X$ be a K3 surface associated with a sextic curve of type $T$. There is the tuple $(H^{2}(X,\mathbb{Z}),H,\Delta)$ associated with $X$. Denote by $\Lambda_{X}\coloneqq H^{2}(X,\mathbb{Z})$. We fix a marking $\phi\colon\Lambda_{X}\to\Lambda$ such that $\phi(H)=l\coloneqq e_{1}+f_{1}$. Then the type IV domain $\mathbb{D}_{H}$ associated with $H^{\perp}\subset\Lambda_{X}$ is identified with $\mathbb{D}_{1}$ by $\phi$.
Let $W$ be the Weyl group of the root lattice $L=\oplus_{i=1}^{k}R_{i}$. Then $W$ acts trivially on $A_{L}$. Thus the action of $W$ on $L$ extends to $\Lambda_{X}$ such that the action on $L^{\perp}$ is identity. It defines a natural inclusion $W\subset\mathrm{O}(\Lambda_{X},H)$. The $W$-action on $\Lambda_{X}$ induces an action on $\Lambda$ by the conjugation of $\phi$, which is by definition $\phi$-equivariant. The group $\mathrm{O}(\Lambda_{X},H)$ is also identified with $\widetilde{\Gamma}_{1}$ via $\mathrm{O}(\Lambda_{X},H)=\phi^{-1}\widetilde{\Gamma}_{1}\phi$.
The period domain $\mathbb{D}_{T}$ is defined to be the type IV domain associated with $P^{\perp}$, where $P^{\perp}$ is the $W$-invariant part of $H^{\perp}$. Namely, $\mathbb{D}_{T}$ is the type IV domain associated with the character subspace of $H^{\perp}\otimes\mathbb{C}$ with respect to the trivial character of $W$.
Define $\widetilde{\Gamma}_{P^{\perp}}\coloneqq\{g\in\mathrm{O}(\Lambda_{X},H)\,\big{|}\,g(P^{\perp})=P^{\perp}\}$ and $\Gamma_{P^{\perp}}\coloneqq\{g\in\mathrm{O}(\Lambda_{X},H)\,\big{|}\,g(\mathbb{D}_{T})=\mathbb{D}_{T}\}$. Note that $[\widetilde{\Gamma}_{T}:\Gamma_{T}]\leq 2$. Let $\widetilde{W}\subset\mathrm{O}(\Lambda_{X},H)$ be the group generated by $W$ and $-\mathrm{Id}$.
Lemma  4.3.
The group $\widetilde{W}$ is the stabilizer of a generic point of $\mathbb{D}_{T}$ in $\mathrm{O}(\Lambda_{X},H)$.
Proof.
Let $x\in\mathbb{D}_{T}$ be a point such that $x^{\perp}\cap\Lambda_{X}=P$. The stabilizer of $x$ in $\mathrm{O}(\Lambda_{X},H)$, denoted by $\mathrm{Stab}(x)$, is contained in $\Gamma_{P^{\perp}}$. Consider the map $\Gamma_{P^{\perp}}\to\mathrm{O}(P^{\perp})$ defined by restriction. Denote by $K$ its kernel, and let $\widetilde{K}\subset\Gamma_{P^{\perp}}$ be the group generated by $K$ and $-\mathrm{Id}$. We claim that $\mathrm{Stab}(x)=\widetilde{K}$. It is clear that $\widetilde{K}\subset\mathrm{Stab}(x)$. For any $g\in\mathrm{O}(\Lambda_{X},H)$ such that $g\cdot v_{x}=\lambda v_{x}$ for some $\lambda\in\mathbb{C}^{*}$, where $v_{x}\in P^{\perp}\otimes\mathbb{C}$ satisfies $[v_{x}]=x\in\mathbb{D}_{T}$. For any $v\in P^{\perp}$, we have
$\langle v,v_{x}\rangle=\langle g\cdot v,g\cdot v_{x}\rangle=\langle g\cdot v,\lambda v_{x}\rangle=\langle\lambda g\cdot v,v_{x}\rangle,$
then by $v_{x}^{\perp}\cap\Lambda_{X}=P$, $g\cdot v=\lambda^{-1}v$. Since $\lambda=\pm 1$, we have $\mathrm{Stab}(x)=\widetilde{K}$.
The Weyl group $W$ extends on $P^{\perp}$ by trivial action, hence $W\subset K$. For any $g\in K$, $g$ acts on $A_{P}$ trivially. Thus $g$ acts on $A_{L}$ trivially. From [Bou02, PLATE I–X], $\mathrm{Aut}(R)/W(R)$ acts faithfully on $A_{R}$ for any irreducible finitely generated positive definite root lattice $R$. Hence $g|_{L}\in\mathrm{O}(L)$ lies in $W$. Thus $K\subset W$. It follows that $\widetilde{K}=\widetilde{W}$. ∎
Let $\widetilde{\Gamma}_{W}$ (respectively, $\Gamma_{W}$) be the normalizer of $W$ in $\mathrm{O}(\Lambda_{X},H)$ (respectively, $\phi^{-1}\Gamma_{1}\phi$). Notice that $\widetilde{\Gamma}_{W}$ (respectively, $\Gamma_{W}$) is also the normalizer of $\widetilde{W}$ in $\mathrm{O}(\Lambda_{X},H)$ (respectively, $\phi^{-1}\Gamma_{1}\phi$). Note that $[\widetilde{\Gamma}_{W}:\Gamma_{W}]\leq 2$. By definition, $\widetilde{\Gamma}_{W}$ preserves the $W$-invariant part of $\Lambda_{X}$, which is $P^{\perp}$.
Lemma  4.4.
We have $\Gamma_{P^{\perp}}=\Gamma_{W}$.
Proof.
We first show that $\widetilde{\Gamma}_{P^{\perp}}=\widetilde{\Gamma}_{W}$. It is clear that $\widetilde{\Gamma}_{P^{\perp}}\supset\widetilde{\Gamma}_{W}$. For any $g\in\mathrm{O}(\Lambda_{X},H)$ such that $g(P^{\perp})=P^{\perp}$, we have $g(L)=L$, namely $g|_{L}\in\mathrm{O}(L)$. Since $W$ is a normal subgroup of $\mathrm{O}(L)$, we have $g^{-1}Wg=W$. Thus $\widetilde{\Gamma}_{P^{\perp}}\subset\widetilde{\Gamma}_{W}$. For any $g\in\widetilde{\Gamma}_{W}$, $\phi g\phi^{-1}$ preserves $\mathbb{D}_{1}$ is equivalent to that $g$ preserves $\mathbb{D}_{T}$. It follows $\Gamma_{P^{\perp}}=\Gamma_{W}$. ∎
Remark  4.5.
Due to [YZ20, Lemma A.4], Lemma 4.4 is a corollary of Lemma 4.3.
Recall that $\widetilde{\Gamma}_{T}$ denotes the image of $\mathrm{O}(\Lambda_{X},\Delta,H)\to\mathrm{O}(P^{\perp})$, and $\Gamma_{T}$ is the subgroup of $\widetilde{\Gamma}_{T}$ that preserves $\mathbb{D}_{T}$ (thus $[\widetilde{\Gamma}_{T}:\Gamma_{T}]\leq 2$). The following result gives an alternative description of the arithmetic group $\Gamma_{T}$.
Lemma  4.6.
The arithmetic group $\Gamma_{W}|_{P^{\perp}}$ coincides with $\Gamma_{T}$.
Proof.
We only need to verify $\widetilde{\Gamma}_{W}|_{P^{\perp}}=\widetilde{\Gamma}_{T}$, namely
$\mathrm{Im}(\{g\in\mathrm{O}(\Lambda_{X},H)\big{|}g^{-1}Wg=W\}\to\mathrm{O}(P^{\perp}))=\mathrm{Im}(\mathrm{O}(\Lambda_{X},\Delta,H)\to\mathrm{O}(P^{\perp})).$
We first verify $\widetilde{\Gamma}_{W}|_{P^{\perp}}\supset\widetilde{\Gamma}_{T}$. An element in $\widetilde{\Gamma}_{T}$ can be expressed as $g|_{P^{\perp}}$ for some $g\in\mathrm{O}(\Lambda_{X},\Delta,H)$. The Weyl group $W$ is generated by the set of reflections $\{s_{\alpha}\,|\,\alpha\in\Delta\}$. For each $\alpha\in\Delta$, we have $g^{-1}s_{\alpha}g=s_{g^{-1}\alpha}$. By definition of $g$, $g^{-1}\alpha$ is also an element in $\Delta$, hence $g^{-1}s_{\alpha}g\in W$. Thus we have $\widetilde{\Gamma}_{W}|_{P^{\perp}}\supset\widetilde{\Gamma}_{T}$.
We next verify $\widetilde{\Gamma}_{W}|_{P^{\perp}}\subset\widetilde{\Gamma}_{T}$. An element $g$ in $\mathrm{O}(\Lambda_{X},H)$ satisfying $g^{-1}Wg=W$ preserves $P$ and $P^{\perp}$. Since $g$ preserves $H$, $g$ also preserves $L$, namely, $g|_{L}\in\mathrm{O}(L)$. Then $g(\Delta)$ is also a base of $L$. The Weyl group $W$ acts on the set of bases of $L$ transitively. Hence there exists a $\sigma\in W$ such that $\sigma\cdot g(\Delta)=\Delta$, namely, $(\sigma g)(\Delta)=\Delta$. By definition, $\sigma g\in\mathrm{O}(\Lambda_{X},\Delta,H)$. Since $W$ acts trivially on $P^{\perp}$, we have $g|_{P^{\perp}}=(\sigma g)|_{P^{\perp}}$. This shows $\widetilde{\Gamma}_{W}|_{P^{\perp}}\subset\widetilde{\Gamma}_{T}$. ∎
By [YZ20, Proposition A.5], we have:
Proposition  4.7.
The natural map $\pi_{T}\colon\Gamma_{T}\backslash\mathbb{D}_{T}\to\Gamma_{1}\backslash\mathbb{D}_{1}$ is a normalization of its image.
Define $\mathcal{H}_{T}^{*}\coloneqq\mathcal{H}_{\infty}\cap\mathbb{D}_{T}$. We will show that $\mathcal{H}_{T}^{*}$ is a $\Gamma_{T}$-invariant hyperplane arrangement in $\mathbb{D}_{T}$, then describe the Looijenga compactification of $\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T}^{*})$.
Proposition  4.8.
We have $\mathcal{H}_{T}^{*}\subset\mathcal{H}_{T}$. In particular, $\mathbb{D}_{T}\not\subset\mathcal{H}_{\infty}$.
Lemma  4.9.
Any root in $L$ has divisibility $1$ in $H^{\perp}$.
Proof.
Assume that there exists a root $r\in L$ with divisibility $2$ in $H^{\perp}$. We will show that $u\coloneqq{1\over 2}(H+r)\in P$. Since $u^{2}=0$, $u\cdot H=1$, this contradicts the fact that $(\Lambda_{X},H,\Delta)$ is an abstract singular type.
We have $H^{\perp}\cong A_{1}\oplus U^{\oplus 2}\oplus E_{8}^{\oplus 3}$ via $\phi$. Let $x\coloneqq\phi^{-1}(e_{1}-f_{1})\in\Lambda_{X}$, then ${1\over 2}(H+x)=\phi^{-1}(e_{1})\in\Lambda_{X}$. The set of roots in $A_{1}\oplus U^{\oplus 2}\oplus E_{8}^{\oplus 3}$ forms two $\mathrm{O}(H^{\perp})$-orbits, which exactly corresponds to divisibility $1$ and $2$. Since the divisibility of $x$ is $2$, there exists $g\in\mathrm{O}(H^{\perp})$ such that $g\cdot x=r$. Since $A_{H^{\perp}}\cong\mathbb{Z}/2$, $g$ acts trivially on $A_{H^{\perp}}$. Thus $g$ extends to an element in $\mathrm{O}(\Lambda_{X},H)$, still denoted by $g$. Hence $u=g\cdot{1\over 2}(H+x)$ lies in $\Lambda_{X}$, then in $P$. ∎
Proof of Proposition 4.8.
For any $\omega\in\mathcal{H}_{T}^{*}$, by the definition of $\mathcal{H}_{\infty}$, there exists a root with divisibility $2$ in $\langle\omega,H\rangle^{\perp}$. By Lemma 4.9, the root does not lie in $L$. Hence $\mathcal{H}_{T}^{*}\subset\mathcal{H}_{T}$. ∎
By Lemma 4.6, $\pi_{T}$ have a natural restriction $\pi_{T}\colon\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T}^{*})\to\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\infty})$. By [YZ20, Theorem A.13] and Lemma 4.3, we have
Proposition  4.10.
There is a natural extension of $\pi_{T}$ to Looijenga compactifications
$\pi_{T}\colon\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathcal{H}_{T}^{*}}\to\overline{\Gamma_{1}\backslash\mathbb{D}_{1}}^{\mathcal{H}_{\infty}},$
which is the normalization of its image.
4.2.3. Identification between Compactifications
Proposition  4.11.
The map $\pi_{T}\colon\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})\to\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\infty})$ is injective. We have the following commutative diagram:
[…]
where the horizontal maps are both isomorphism and the vertical maps are both injective.
Proof.
By Lemma 4.4 and Lemma 4.6, we have $\Gamma_{T}=\{g|_{P^{\perp}}\,|\,g\in\Gamma_{1},g(P^{\perp})=P^{\perp}\}$. Namely, for any $\omega,\omega^{\prime}\in\mathbb{D}_{T}\subset\mathbb{D}_{1}$, if there exists a $g\in\Gamma_{1}$ such that $g\cdot\omega=\omega^{\prime}$, then $g|_{P^{\perp}}\in\Gamma_{T}$. Hence $\pi_{T}\colon\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})\to\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\infty})$ is injective.
The period map $\mathscr{P}\colon\mathcal{M}^{\mathrm{ADE}}\xrightarrow[]{\cong}\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\infty})$ is compatible with the period map of marked complex K3 surfaces. By Lemma 3.8, $\mathscr{P}_{T}\colon\mathcal{M}_{T}\xrightarrow[]{\cong}\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$ is also compatible with the period map of marked complex K3 surfaces. By Theorem 3.4, Theorem 4.1, together with $j_{T}\colon\mathcal{M}_{T}\hookrightarrow\mathcal{M}^{\mathrm{ADE}}\hookrightarrow\overline{\mathcal{M}}$ and $\pi_{T}\colon\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})\hookrightarrow\Gamma_{1}\backslash(\mathbb{D}_{1}-\mathcal{H}_{\infty})\hookrightarrow\overline{\Gamma_{1}\backslash\mathbb{D}_{1}}^{\mathcal{H}_{\infty}}$, we have the claimed commutative diagram. ∎
Theorem  4.12.
The period map $\mathscr{P}_{T}\colon\mathcal{M}_{T}\xrightarrow[]{\cong}\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$ extends to an isomorphism between the GIT compactification and the Looijenga compactification $\mathscr{P}_{T}\colon\overline{\mathcal{M}}_{T}\to\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathcal{H}_{T}^{*}}$. We have the following commutative diagram:
[…]
where the two vertical maps are normalizatons of their images.
Proof.
By Proposition 4.11, $j_{T}(\mathcal{M}_{T})$ and $\pi_{T}(\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T}))$ are isomorphic via $\mathscr{P}$. Then their (Zariski) closures in $\overline{\mathcal{M}}$ and $\overline{\Gamma_{1}\backslash\mathbb{D}_{1}}^{\mathcal{H}_{\infty}}$ are isomorphic via $\mathscr{P}$. Thus the normalizations of the two closures, namely ¯$\overline{\mathcal{M}}_{T}$ and $\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathcal{H}_{T}^{*}}$, are isomorphic via a unique extension of $\mathscr{P}_{T}\colon\mathcal{M}_{T}\xrightarrow[]{\cong}\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$, which is still denoted by $\mathscr{P}_{T}$.
By Proposition 4.2 and Proposition 4.10, $j_{T}$ and $\pi_{T}$ extend to the compactifications ¯$\overline{\mathcal{M}}_{T}$ and $\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathcal{H}_{T}}$. By Proposition 4.11, the compactified diagram is also commutative. ∎
5. More Explicit Characterization of Lattices
In this section, we consider a plane curve of even degree. For a singular type $T$, let $F$ be an element in $\mathcal{V}_{T}$. We omit the index $F$ of all relevant symbols in this section for simplicity. Our main goal is to give an explicit description of the lattice $P=P_{F}$.
5.1. A Topological Lemma
We first state a general lemma for branched double covers of smooth manifolds. Let $p\colon X\to S$ be a double cover and the branched locus be the disjoint union of submanifolds $C_{1},\cdots,C_{l}$ of $S$ of real codimension two. Then $X$ and $S$ are smooth. Denote by $\iota$ the deck transformation of $p$. Let $H^{2}(X,\mathbb{Z})^{\iota}$ be the fixed sublattice of the induced action of $\iota$ on $H^{2}(X,\mathbb{Z})$.
Lemma  5.1.
The double cover $p$ induces a lattice saturation $H^{2}(S,\mathbb{Z})(2)\hookrightarrow H^{2}(X,\mathbb{Z})^{\iota}$ such that
$\frac{H^{2}(X,\mathbb{Z})^{\iota}}{H^{2}(S,\mathbb{Z})(2)}\cong(\mathbb{Z}/2)^{l-1}.$
Proof.
This lemma is a slight generalization of [YZ23, Prop $3.4$] and the proof is similar. We provide the proof here for completeness.
Let $M$ be the lattice generated by $H^{2}(S,\mathbb{Z})(2)$ and $[\pi^{-1}C_{1}],\cdots,[\pi^{-1}C_{l}]$. By [YZ23, Lemma $3.3$], we have $M\subset H^{2}(X,\mathbb{Z})^{\iota}$ with the same rank, denoted by $r$. Note that the rank of $H^{2}(S,\mathbb{Z})$ also equals $r$. We have $A_{H^{2}(S,\mathbb{Z})(2)}\cong(\mathbb{Z}/2)^{r}$ because $H^{2}(S,\mathbb{Z})$ is unimodular. By the consecutive inclusions $H^{2}(S,\mathbb{Z})(2)\subset M\subset H^{2}(X,\mathbb{Z})^{\iota}\subset(H^{2}(S,\mathbb{Z})(2))^{\vee}$, the quotients $H^{2}(X,\mathbb{Z})^{\iota}/H^{2}(S,\mathbb{Z})(2)$ and $M/H^{2}(S,\mathbb{Z})(2)$ are isomorphic to certain powers of $\mathbb{Z}/2$. Tensoring the short exact sequence
$0\to H^{2}(S,\mathbb{Z})\to H^{2}(X,\mathbb{Z})\to H^{2}(X,\mathbb{Z})/H^{2}(S,\mathbb{Z})(2)\to 0$
with $\mathbb{Z}/2$, we obtain a long exact sequence
00 0 $\to\mathrm{Tor}(H^{2}(X,\mathbb{Z})/H^{2}(S,\mathbb{Z})(2),\mathbb{Z}/2)\to H^{2}(S,\mathbb{Z}/2)\to H^{2}(X,\mathbb{Z}/2)$
$\to(H^{2}(X,\mathbb{Z})/H^{2}(S,\mathbb{Z})(2))\otimes\mathbb{Z}/2\to 0.$
The $\mathbb{Z}/2$-rank of the kernel of $H^{2}(S,\mathbb{Z}/2)\to H^{2}(X,\mathbb{Z}/2)$ equals to the $\mathbb{Z}/2$-rank of the $2$-torsion part (in this case also the whole torsion part) $(H^{2}(X,\mathbb{Z})/H^{2}(S,\mathbb{Z})(2))_{tor}$ of
$H^{2}(X,\mathbb{Z})/H^{2}(S,\mathbb{Z})(2)$. By [LW95, Theorem 1], we have the exact sequence
$0\to H^{1}(S,\bigcup_{i=1}^{l}C_{l},\mathbb{Z}/2)\to H^{2}(S,\mathbb{Z}/2)\to H^{2}(X,\mathbb{Z}/2).$
There is a long exact sequence for the relative cohomology groups
$0\to H^{0}(S,\mathbb{Z}/2)\to H^{0}(\bigcup_{i=1}^{l}C_{i},\mathbb{Z}/2)\to H^{1}(S,\bigcup_{i=1}^{l}C_{i},\mathbb{Z}/2)\to 0$
It follows that the $\mathbb{Z}/2$-rank of $H^{1}(S,\bigcup_{i=1}^{l}C_{i},\mathbb{Z}/2)$ is $l-1$. From the above exact sequences we conclude that the $\mathbb{Z}/2$-rank of $(H^{2}(X,\mathbb{Z})/H^{2}(S,\mathbb{Z})(2))_{tor}$ is $l-1$. Since $H^{2}(X,\mathbb{Z})^{\iota}$ is the saturation of $H^{2}(S,\mathbb{Z})$ in $H^{2}(X,\mathbb{Z})$, we have $H^{2}(X,\mathbb{Z})^{\iota}/H^{2}(S,\mathbb{Z})(2)\cong(\mathbb{Z}/2)^{l-1}$. ∎
5.2. Involution of Root Lattice
Let $Z=Z(F)$ be a plane curve of even degree. Suppose $Z$ has only simple singularities. Let $\widehat{X}$ denote the double cover of $\mathbb{P}^{2}$ branched along $Z$, and $X$ denote the minimal resolution of $\widehat{X}$. Similar to the sextic case, there is a regular involution $\iota$ of $X$, with quotient $S=X/\iota$. Denote by $f\colon S\to\mathbb{P}^{2}$ the natural birational morphism, which is the composition of consecutive blowups. The strict transform $\widetilde{Z}\subset S$ of $Z$ is smooth. The double cover $X\to S$ is branched along the disjoint union of $\widetilde{Z}$ and some exceptional curves, see §2.2. We have $H^{2}(S,\mathbb{Z})\cong\left\langle f^{*}h\right\rangle\oplus T$, where $T$ arises from the blowups of $S\to\mathbb{P}^{2}$.
For a root lattice of type $R$, we denote by $w_{0}(R)$ the longest element of the Weyl group $W(R)$ with respect to a given base. Recall that the root lattice $L=L_{F}$ is the direct sum of all $L(R)$, where $R$ runs through all singularity types (with multiplicities) appearing in $\widehat{X}$. The longest element $w_{0}(L)\in W(L)$ is equal to the product of all $w_{0}(R)$. We have $-w_{0}(L)\in\mathrm{O}(L)$. For simplicity, the induced $\iota$-action on $H^{2}(X,\mathbb{Z})$ (and its restrictions on subspaces) is still denoted by $\iota$. Denote by $L^{\iota}$ the sublattice fixed by $\iota$.
Proposition  5.2.
We have $\iota=-w_{0}(L)$ as automorphisms of $L$.
Proof.
It suffices to check this for every irreducible root lattice $L$. Notice that $-w_{0}(L)$ is an involution preserving the base $\Delta$. The graph automorphisms induced by $w_{0}(L)$ are listed in [Bou02, PLATE I–X]. ∎
5.3. Comparison of Lattices
The double cover $p\colon X\to S$ induces an injective morphism $p^{*}\colon H^{2}(S,\mathbb{Z})(2)\to H^{2}(X,\mathbb{Z})$ of lattices, which maps $f^{*}h$ to $H$. It induces a saturation $T(2)\hookrightarrow L^{\iota}$.
Lemma  5.3.
We have $L^{\iota}/T(2)\cong(\mathbb{Z}/2)^{m}$, where $m$ is the number of exceptional curves in $S$ lying in the branch locus of $X\to S$.
Proof.
It suffices to check this for every irreducible root lattice $L$. Suppose $D$ is a branch exceptional curve in $S$, then the Chern class of $\pi^{-1}(D)$ lies in $L^{\iota}$, while twice of it lies in $T(2)$. This implies the required isomorphism. ∎
Table 1 provides the values of $m$ for root lattices of ADE types.
Table 1.
Root Type $R$ $A_{n}$ $D_{n}$ $E_{6}$ $E_{7}$ $E_{8}$
$m_{R}$ 00 $[\frac{n-2}{2}]$ $1$ $3$ $4$
Corollary  5.4.
If the plane sextic curve $Z$ has only singularities of type $A_{1},D_{2n}(n\geq 2),E_{7},E_{8}$, then $L^{\iota}=L$. If only type $A_{1}$ appears, then $T(2)\cong L$ as lattices.
For a root lattice $R$ of ADE type, we give a description of $R^{w_{0}(R)}$, which is the anti-invariant part of the $-w_{0}(R)$-action. Denote by $r_{R}^{-}$ the rank of $R^{w_{0}(R)}$.
Lemma  5.5.
The discriminant group of $R^{w_{0}(R)}$ for each ADE-type $R$ is listed in Table 2.
Table 2.
Root Type $R$ $A_{2n-1}$ $A_{2n}$ $D_{n}$ (odd) $E_{6}$
$r_{R}^{-}$ $n-1$ $n$ $1$ $2$
$A_{R^{w_{0}(R)}}$ $(\mathbb{Z}/2)^{n-2}\times\mathbb{Z}/2n$ $(\mathbb{Z}/2)^{n}\times\mathbb{Z}/(2n+1)$ $\mathbb{Z}/4$ $\mathbb{Z}/3\times(\mathbb{Z}/2)^{2}$
For the type $A_{1}$, $D_{n}$ ($n$ even), $E_{7}$ and $E_{8}$, $R^{w_{0}(R)}$ is trivial.
Denote by $D_{1},\cdots,D_{l^{\prime}}$ the irreducible components of $Z\subset\mathbb{P}^{2}$. Let $\widetilde{D}_{1},\cdots,\widetilde{D}_{l^{\prime}}$ be the connected components of the strict transform of $\widetilde{Z}\subset S$. Denote by $E_{1},\cdots,E_{m}$ the exceptional curves of $S\to\mathbb{P}^{2}$ lying in the branched locus of $\pi\colon X\to S$. The double cover $\pi$ is branched along the disjoint union of $\widetilde{D}_{1},\cdots,\widetilde{D}_{l^{\prime}}$ and $E_{1},\cdots,E_{m}$. Let $l\coloneqq l^{\prime}+m$.
Recall $P=P_{F}$ is the primitive hull of $\left\langle H\right\rangle\oplus L$ in $H^{2}(X,\mathbb{Z})$. Let $M$ be the sublattice of $H^{2}(X,\mathbb{Z})$ generated by $H$, $L$ and Chern classes of curves $\pi^{-1}(\widetilde{D}_{1}),\cdots,\pi^{-1}(\widetilde{D}_{l^{\prime}})\subset X$. Denote by $\beta_{i}$ the Chern class of $\pi^{-1}(\widetilde{D}_{i})\subset X$.
Lemma  5.6.
We have
$\frac{P^{\iota}}{(\left\langle H\right\rangle\oplus L)^{\iota}}\cong(\mathbb{Z}/2)^{l^{\prime}-1}.$
Proof.
By definition $\beta_{i}\in H^{2}(X,\mathbb{Z})^{\iota}$ for $1\leq i\leq l^{\prime}$. Denote by $N$ the lattice $H^{2}(S,\mathbb{Z})(2)=(\left\langle f^{*}h\right\rangle\oplus T)(2)$, which is a sublattice of $H^{2}(X,\mathbb{Z})$. By Lemma 5.1, $H^{2}(X,\mathbb{Z})^{\iota}$ is a saturation of $N$ in $H^{2}(X,\mathbb{Z})$. Since $N\subset\left\langle H\right\rangle\oplus L$, we have $H^{2}(X,\mathbb{Z})^{\iota}\subset P$. It follows $H^{2}(X,\mathbb{Z})^{\iota}=P^{\iota}$. Consequently, we have $M\subset P$. Then there are successive saturations of lattices
$N\hookrightarrow(\left\langle H\right\rangle\oplus L)^{\iota}\hookrightarrow M^{\iota}\hookrightarrow P^{\iota}.$ (3)
By Lemma 5.1, we have $P^{\iota}/N\cong(\mathbb{Z}/2)^{l-1}$. By Lemma 5.3, we have $(\left\langle H\right\rangle\oplus L)^{\iota}/N\cong(\mathbb{Z}/2)^{m}$. Thus the lemma follows. ∎
Lemma  5.7.
Any $l^{\prime}-1$ elements among $\beta_{1},\cdots,\beta_{l^{\prime}}$ are $\mathbb{Z}/2$-independent in $\frac{M^{\iota}}{(\left\langle H\right\rangle\oplus L)^{\iota}}$.
Proof.
The nodal case is proved in [YZ23, Proposition 3.4]. Without loss of generality, we consider the first $l^{\prime}-1$ elements $\beta_{1},\cdots,\beta_{l^{\prime}-1}$. Consider the set $A=(\mathop{\bigcup}\limits_{i=1}^{l^{\prime}-1}D_{i})\cap D_{l^{\prime}}$. The singularities in $A$ must be of type $A_{2n-1},D_{n},E_{7}$. It follows from the condition that each singularity in $A$ has at least two local components.
It suffices to show that any nontrivial linear combination of $\beta_{1},\cdots,\beta_{l^{\prime}-1}$ with coefficients 00 or $1$ does not belong to $\langle H\rangle\oplus L$. It is enough to consider $\Sigma=\beta_{1}+\cdots+\beta_{k}$, where $k$ is an integer satisfying $11\leq k\leq l^{\prime}-$ 1. Suppose there is a singularity of type $A_{2n-1}$ on $(\mathop{\bigcup}\limits_{i=1}^{k}D_{i})\cap D_{l^{\prime}}$. The singularity lies on exactly two irreducible components of the divisor, which we denote by $D_{i}$ and $D_{l^{\prime}}$. The graph of exceptional curves and strict transforms of $D_{i}$, $D_{l^{\prime}}$ is drawn in Figure 1 (also see [BHPVdV04, Chapter III, §7, Table 1]).
Figure 1. Resolution graph of $A_{2n-1}$-type
Both $\beta_{i}$ and $\beta_{l^{\prime}}$ intersect only the middle curve. Hence the projection of $\beta_{i}$ (also $\beta_{l^{\prime}}$) to the root lattice of type $A_{2n-1}$ is the fundamental weight $\lambda_{n}$. The weight $\lambda_{n}$ can be expressed as
$\lambda_{n}=-{1\over 2}x_{1}-{2\over 2}x_{2}-{3\over 2}x_{3}-\cdots-{n\over 2}x_{n}-{n-1\over 2}x_{n+1}-\cdots-{1\over 2}x_{2n-1},$
where the coefficients of both $x_{i}$ and $x_{2n-i}$ $(1\leq i\leq n)$ are $-{i\over 2}$. (A general calculation of fundamental weights can be found in [Bou02, PLATE I–X].) This is exactly the projection of $\Sigma$ to the associated $A_{2n-1}$-root lattice. Thus $\Sigma\notin\langle H\rangle\oplus L$.
The cases of types $D_{n}$ and $E_{7}$ are similar, but a bit more involved. We call the exceptional curve corresponding to $m$-node the $m$-line. See Figure 2, 3, 4 for the numbering of each node.
Figure 2. Resolution graph of $D_{2n-1}$-type
Figure 3. Resolution graph of $D_{2n}$-type
Figure 4. Resolution graph of $E_{7}$-type
A singularity of type $D_{2n-1}$ or $E_{7}$ lies on exactly two irreducible components, which we denote by $D_{j}$ and $D_{l^{\prime}}$. In $D_{2n-1}$ case, the strict transform of the locally cuspidal component intersects only the $(2n-3)$-line. The strict transform of the locally smooth component intersects only the $1$-line. The projection of $\beta_{i}$ to the root lattice of type $D_{2n-1}$ is the fundamental weight $\lambda_{1}$ or $\lambda_{2n-3}$. The weights $\lambda_{1}$ and $\lambda_{2n-3}$ have the following expressions
$\lambda_{2n-3}$ $=-x_{1}-2x_{2}-3x_{3}-\cdots-(2n-3)x_{2n-3}-{2n-3\over 2}(x_{2n-2}+x_{2n-1}),$
$\lambda_{1}$ $=-(x_{1}+x_{2}+\cdots+x_{2n-3})-{1\over 2}(x_{2n-2}+x_{2n-1}).$
In $E_{7}$ case, the strict transform of the locally cuspidal component intersects only the $2$-line. The strict transform of the locally smooth component intersects only the $7$-line. The weights $\lambda_{2}$ and $\lambda_{7}$ have the following expressions
$\lambda_{2}$ $=-{1\over 2}(4x_{1}+7x_{2}+8x_{3}+12x_{4}+9x_{5}+8x_{6}+3x_{7}),$
$\lambda_{7}$ $=-{1\over 2}(2x_{1}+3x_{2}+4x_{3}+6x_{4}+5x_{5}+4x_{6}+3x_{7}).$
These are exactly the projections of $\Sigma$ to the associated $D_{2n-1}$ and $E_{7}$. Thus $\Sigma\notin\langle H\rangle\oplus L$.
A singularity of type $D_{2n}$ lies on exactly three irreducible components, which we denote by $D_{k}$, $D_{l}$ and $D_{l^{\prime}}$. The strict transform of two components intersect only with the $2n$-line and the $(2n-1)$-line respectively. The strict transform of the rest component intersects only the $1$-line. The corresponding fundamental weights can be expressed as
$\lambda_{1}$ $=-(x_{1}+x_{2}+\cdots+x_{2n-2})-{1\over 2}(x_{2n-1}+x_{2n}),$
$\lambda_{2n-1}$ $=-{1\over 2}x_{1}-{2\over 2}x_{2}-{3\over 2}x_{3}-\cdots-{2n-2\over 2}x_{2n-2}-{n\over 2}x_{2n-1}-{n-1\over 2}x_{2n},$
$\lambda_{2n}$ $=-{1\over 2}x_{1}-{2\over 2}x_{2}-{3\over 2}x_{3}-\cdots-{2n-2\over 2}x_{2n-2}-{n-1\over 2}x_{2n-1}-{n\over 2}x_{2n}.$
The strict transform of $D_{l^{\prime}}$ may give one or two components in the resolution graph of the singularity. Every linear combination of the rest components with coefficients 00 or $1$ has non-integer coefficients. These are exactly the projections of $\Sigma$ to the associated $D_{2n}$. Thus $\Sigma\notin\langle H\rangle\oplus L$.
By the above calculation, any nontrivial $\mathbb{Z}/2$-combination of $\beta_{1},\cdots\beta_{l^{\prime}-1}$ does not lie in $\langle H\rangle\oplus L$. Hence $\beta_{1},\cdots\beta_{l^{\prime}-1}$ is $\mathbb{Z}/2$-independent in $\frac{M^{\iota}}{(\left\langle H\right\rangle\oplus L)^{\iota}}$. ∎
Remark  5.8.
The above calculation in the case of sextic curves can also be found in [Yan96, §3, Tabel 1].
Proposition  5.9.
We have
$\frac{M}{\left\langle H\right\rangle\oplus L}\cong(\mathbb{Z}/2)^{l^{\prime}-1}.$
This finite group is generated by equivalence classes of $\beta_{1},\cdots,\beta_{l^{\prime}-1}$. Moreover, $M^{\iota}=P^{\iota}$.
Proof.
Since $\beta_{i}\in M^{\iota}$ for $1\leq i\leq l^{\prime}$, the saturations in (3) induces injective maps of the following finite abelian groups
$\frac{M^{\iota}}{(\left\langle H\right\rangle\oplus L)^{\iota}}\hookrightarrow\frac{P^{\iota}}{(\left\langle H\right\rangle\oplus L)^{\iota}}\hookrightarrow\frac{P}{\left\langle H\right\rangle\oplus L}.$
By Lemma 5.6, we have
$\frac{M^{\iota}}{(\left\langle H\right\rangle\oplus L)^{\iota}}=\frac{P^{\iota}}{(\left\langle H\right\rangle\oplus L)^{\iota}}\cong(\mathbb{Z}/2)^{l^{\prime}-1}.$
By Lemma 5.7, there is only one nontrivial relation among $\beta_{1},\cdots,\beta_{l^{\prime}}$. This isomorphism implies that $M^{\iota}=P^{\iota}$. Moreover, the inclusion
$\frac{M^{\iota}}{(\left\langle H\right\rangle\oplus L)^{\iota}}\hookrightarrow\frac{M}{\left\langle H\right\rangle\oplus L}$
is surjective (hence an isomorphism) because $\beta_{1},\cdots,\beta_{l^{\prime}}$ lies in $M^{\iota}$. We have done. ∎
Lemma  5.10.
Let $L$ be an even lattice and $\iota\in\mathrm{O}(L)$ be an involution. Then $\frac{L}{L^{\iota}\oplus L^{-\iota}}$ is $2$-elementary, and its cardinality satisfies
$\bigg{|}\frac{L}{L^{\iota}\oplus L^{-\iota}}\bigg{|}\leq 2^{min\{\mathrm{rank}(L^{\iota}),\mathrm{rank}(L^{-\iota})\}}.$
In particular, for every root lattice $L$ of ADE type, the equality is achieved, i.e.
$\frac{L}{L^{\iota}\oplus L^{-\iota}}\cong(\mathbb{Z}/2)^{\mathrm{rank}(L^{-\iota})}.$
Proof.
For any $v\in L$, we have $2v=(v+\iota(v))+(v-\iota(v))$, hence $2v\in L^{\iota}\oplus L^{-\iota}$. Thus $\frac{L}{L^{\iota}\oplus L^{-\iota}}$ is $2$-elementary. Denote by $r\coloneqq min\{\mathrm{rank}(L^{\iota}),\mathrm{rank}(L^{-\iota})\}$. For any $n>r$, we assume that there exist $\frac{v_{1}+w_{1}}{2},\cdots,\frac{v_{n}+w_{n}}{2}$ in $L\backslash(L^{\iota}\oplus L^{-\iota})$. Without loss of generality, we assume that $\mathrm{rank}(L^{\iota})\geq\mathrm{rank}(L^{-\iota})$. Since $n>r=\mathrm{rank}(L^{-\iota})$, there exist $a_{1},\cdots,a_{n}$ ($\exists a_{i}\neq 0$) such that $\sum\limits_{i=1}^{n}a_{i}w_{i}=0$. Then $\sum\limits_{i=1}^{n}a_{i}\frac{v_{i}+w_{i}}{2}=\sum\limits_{i=1}^{n}a_{i}\frac{v_{i}}{2}$ lies in $L^{\iota}$. Hence $\sum\limits_{i=1}^{n}a_{i}\frac{v_{i}+w_{i}}{2}$ vanishes in $\frac{L}{L^{\iota}\oplus L^{-\iota}}$, namely, there exists a nontrivial relation among $\frac{v_{1}+w_{1}}{2},\cdots,\frac{v_{n}+w_{n}}{2}$. Thus $\dim_{\mathbb{Z}/2}(\frac{L}{L^{\iota}\oplus L^{-\iota}})\leq\mathrm{rank}(L^{-\iota})$ (view $\frac{L}{L^{\iota}\oplus L^{-\iota}}$ as a $\mathbb{Z}/2$-vector space).
For root lattices of ADE types, the equality arises from a case-by-case check. Notice that
$\bigg{|}\frac{L}{L^{\iota}\oplus L^{-\iota}}\bigg{|}^{2}=\frac{|A_{L^{\iota}}\oplus A_{L^{-\iota}}|}{|A_{L}|}.$
For each ADE type, $A_{L^{-\iota}}$ is listed in Table 2. We list $A_{L^{\iota}}$ in Table 3.
Table 3.
Root Type $R$ $A_{2n-1}$ $A_{2n}$ $D_{n}$ (odd) $E_{6}$
$r_{R}^{+}$ $n$ $n$ $n-1$ $4$
$A_{R^{-w_{0}(R)}}$ $(\mathbb{Z}/2)^{n}$ $(\mathbb{Z}/2)^{n}$ $(\mathbb{Z}/2)^{2}$ $(\mathbb{Z}/2)^{2}$
Proposition  5.11.
The lattice $M$ is the biggest $\mathbb{Z}/2$-saturation of $P^{\iota}\oplus L^{-\iota}$ in $H^{2}(X,\mathbb{Z})$, and $\frac{M}{P^{\iota}\oplus L^{-\iota}}\cong(\mathbb{Z}/2)^{r^{-}}$.
Proof.
Since $P^{\iota}$ is a saturation of $\left\langle H\right\rangle\oplus L^{\iota}$, $L^{-\iota}$ is orthogonal to $P^{\iota}$ in $P$. Consider the saturation
$P^{\iota}\oplus L^{-\iota}\hookrightarrow P.$
By Proposition 5.9, we have $M^{\iota}=P^{\iota}$. Thus there is a natural saturation
$P^{\iota}\oplus L^{-\iota}\hookrightarrow M.$
We have the following successive saturations
$\left\langle H\right\rangle\oplus L^{\iota}\oplus L^{-\iota}\hookrightarrow\left\langle H\right\rangle\oplus L\hookrightarrow M.$
By Lemma 5.10, we have $\frac{L}{L^{\iota}\oplus L^{-\iota}}\cong(\mathbb{Z}/2)^{r^{-}}$. By Proposition 5.9, we have $\frac{M}{\left\langle H\right\rangle\oplus L}\cong(\mathbb{Z}/2)^{l^{\prime}-1}$. Moreover, we have another successive saturations
$\left\langle H\right\rangle\oplus L^{\iota}\oplus L^{-\iota}\hookrightarrow P^{\iota}\oplus L^{-\iota}\hookrightarrow M.$
By Lemma 5.6 and Proposition 5.9, we have $\frac{P^{\iota}}{\left\langle H\right\rangle\oplus L^{\iota}}=\frac{M}{\left\langle H\right\rangle\oplus L}\cong(\mathbb{Z}/2)^{l^{\prime}-1}$. It is direct to verify that $(\left\langle H\right\rangle\oplus L)\cap(P^{\iota}\oplus L^{-\iota})=\left\langle H\right\rangle\oplus L^{\iota}\oplus L^{-\iota}$. Thus $\frac{M}{\left\langle H\right\rangle\oplus L^{\iota}\oplus L^{-\iota}}\cong(\mathbb{Z}/2)^{l^{\prime}+r^{-}-1}$. Therefore we conclude
$\frac{M}{P^{\iota}\oplus L^{-\iota}}\cong(\mathbb{Z}/2)^{r^{-}},$
where $r^{-}$ is the sum of all $r_{R}^{-}$ with $R$ that appear in $L$ (with multiplicities).
Since $P^{\iota}$ is primitive in $P$, by lattice theory, the quotient $\frac{P}{P^{\iota}\oplus L^{-\iota}}$ is isomorphic to certain subgroup of $A_{L^{-\iota}}$. By Lemma 5.5, the biggest subgroup of each $A_{R^{w_{0}(R)}}$ of the form $(\mathbb{Z}/2)^{k}$ is isomorphic to $(\mathbb{Z}/2)^{r_{R}^{-}}$. Notice that $(\mathbb{Z}/2)^{r^{-}}$ can be naturally decomposed into the product $\prod\limits_{R}(\mathbb{Z}/2)^{l_{T}(R)r_{R}^{-}}$. Therefore $M$ is the biggest possible $\mathbb{Z}/2$-saturation of $P^{\iota}\oplus L^{-\iota}$. ∎
Remark  5.12.
Since $H$ is primitive in $H^{2}(X,\mathbb{Z})$, $\frac{P}{\left\langle H\right\rangle\oplus L}$ is a subgroup of $A_{L}$.
Remark  5.13.
There exist examples such that $M$ is a proper sublattice of $P$, for example, see §7.3 and Propostion 7.16. By [YZ23, Proposition 3.4], $M=P$ holds for all nodal singular types. We will give more examples with $M=P$ in §7.
6. Orbifold Structures
Given a singular type $T$, both the moduli space $\mathcal{M}_{T}$ and the associated arithmetic quotient $\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$ admit natural orbifold structures induced by their nontrivial automorphism groups. It is natural to study the relationship of these two orbifold structures. See [KR12] and [Zhe21] for some well-known cases. In this section, we compare the orbifold structures of the two sides of the period map $\mathscr{P}_{T}\colon\mathcal{M}_{T}\to\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$. We follow the setup in §3.
6.1. A Lattice-theoretic Criterion
For $Z(F)\in\mathbb{P}\mathcal{V}_{T}$, let $(X_{F},H_{F},\iota_{F})$ be defined as in §2.3, and $\Delta_{F}$ be defined as in §3.1.
Proposition  6.1.
We have a natural isomorphism
$\mathrm{Stab}_{Z(F)}$ $\cong\mathrm{Aut}(X_{F},H_{F})/\{\mathrm{Id},\iota_{F}\}$
$\cong\mathrm{O}(\Lambda_{F},H_{F},\Delta_{F},H^{2,0}(X_{F}))/\{\mathrm{Id},\iota_{F}^{*}\},$
where $\mathrm{Stab}_{Z(F)}$ is the stabilizer of $Z(F)$ in $\mathrm{PGL}(3)$.
Lemma  6.2.
We have
$\mathrm{Aut}(X_{F},H_{F},\Delta_{F})$ $=\mathrm{Aut}(X_{F},H_{F})$
$\cong\mathrm{O}(\Lambda_{F},H_{F},\Delta_{F},H^{2,0}(X_{F})).$
Proof.
Let $f\in\mathrm{Aut}(X_{F})$ satisfy $f^{*}H_{F}=H_{F}$. Then $f^{*}L_{F}=L_{F}$, as $L_{F}$ is the root lattice of $\langle H_{F},H^{2,0}(X_{F})\rangle^{\perp}_{\Lambda_{F}}$. Since $f^{*}$ maps effective classes to effective classes, it follows that $f^{*}\Delta_{F}=\Delta_{F}$. Consequently, we conclude $\mathrm{Aut}(X_{F},H_{F},\Delta_{F})=\mathrm{Aut}(X_{F},H_{F})$.
Moreover, by the proof of Theorem 3.3, we can construct an ample class $\alpha_{F}$ of $X_{F}$ lying in $\langle H_{F}\rangle\oplus L_{F}$. Since $\mathrm{Aut}(X_{F},\alpha_{F})\cong\mathrm{O}(\Lambda_{F},\alpha_{F},H^{2,0}(X_{F}))$ (given by $f\mapsto f^{*}$) and $\mathrm{Aut}(X_{F},H_{F},\Delta_{F})\subset\mathrm{Aut}(X_{F},\alpha_{F})$, it follows that $f\mapsto f^{*}$ gives $\mathrm{Aut}(X_{F},H_{F},\Delta_{F})\cong\\ \mathrm{O}(\Lambda_{F},H_{F},\Delta_{F},H^{2,0}(X_{F}))$. ∎
Lemma  6.3.
Each element $f\in\mathrm{O}(\Lambda_{F},\Delta_{F},H_{F})$ commutes with the involution $\iota_{F}$.
Proof.
Consider a singularity of type $R$ on $Z(F)$. The lattices $L(R)$ and $f(L(R))$, both of type $R$, satisfy $f(L(R)\cap\Delta_{F})=f(L(R))\cap\Delta_{F}$. By Proposition 2.6 and Remark 2.7, we can identify the generators of $L(R)^{\iota_{F}}$ and $f(L(R))^{\iota_{F}}$. This implies $f(L(R)^{\iota_{F}})=f(L(R))^{\iota_{F}}$. Since $f(H_{F})=H_{F}$, we conclude $f(\Lambda_{F}^{\iota_{F}})=\Lambda_{F}^{\iota_{F}}$. Consequently, $f$ commutes with $\iota_{F}$. ∎
Lemma  6.4.
Every automorphism $f\in\mathrm{Aut}(X_{F},H_{F})$ naturally induces an automorphism of $Z(F)\subset\mathbb{P}^{2}$. The only automorphisms of $X_{F}$ inducing the identity map on $Z(F)$ are the involution $\iota_{F}$ and the identity map.
Proof.
By Lemma 6.2, every $f\in\mathrm{Aut}(X_{F},H_{F})$ induces a $f^{*}\in\mathrm{O}(\Lambda_{F})$ that preserves $H_{F}$ and $\Delta_{F}$. Lemma 6.3 ensures that $f^{*}$ commutes with $\iota_{F}^{*}$ on $\Lambda_{F}$. By the injectivity of $\mathrm{Aut}(X_{F})\hookrightarrow\mathrm{O}(\Lambda_{F})$, it follows that $f$ commutes with $\iota_{F}$. Hence $f$ induces an automorphism of $S_{F}$ that preserves the strict transform of $Z(F)$, which lies in the branched locus of $p_{F}\colon X_{F}\to S_{F}$. Moreover, since $f^{*}H_{F}=H_{F}$, $f$ descends to an automorphism $\overline{f}$ of $\mathbb{P}^{2}$ that preserves $Z(F)$.
Let $f\in\mathrm{Aut}(X_{F})$ be an automorphism such that the induced map $\overline{f}$ restricts to the identity map on $Z(F)$. Then $\overline{f}$ is the identity map on $\mathbb{P}^{2}$. The identity map lifts trivially to $S_{F}$, and thus $f$ must coincide with either the identity map on $X_{F}$ or the involution $\iota_{F}$. ∎
Proof of Proposition 6.1.
There exists a natural map
$\mathrm{Aut}(X_{F},H_{F})\to\mathrm{Stab}_{Z(F)},\quad f\mapsto\overline{f}.$
By Lemma 6.4, the kernel of this map is $\{\mathrm{Id},\iota_{F}\}$. Moreover, every automorphism of $\mathbb{P}^{2}$ preserving $Z(F)$ lifts to an automorphism of the double cover branched along $Z(F)$, and subsequently to an automorphism of $X_{F}$. Thus we obtain $\mathrm{Stab}_{Z(F)}\cong\mathrm{Aut}(X_{F},H_{F})/\{\mathrm{Id},\iota_{F}\}$.
By Lemma 6.2, there is a natural isomorphism
$\mathrm{Aut}(X_{F},H_{F})/\{\mathrm{Id},\iota_{F}\}\cong\mathrm{O}(\Lambda_{F},H_{F},\Delta_{F},H^{2,0}(X_{F}))/\{\mathrm{Id},\iota_{F}^{*}\}.$
Combining these results, the proposition follows. ∎
Let $Z(F_{0})\in\mathbb{P}\mathcal{V}_{T}$ be a base point. By Corollary 2.14, the involution $\iota_{F_{0}}$ induces an action on $\Lambda_{F_{0}}$ preserving $H_{F_{0}}$ and $\Delta_{F_{0}}$, and acts as $-\mathrm{Id}$ on $Q_{F_{0}}$. We denote by $P\Gamma_{T}$ the quotient group $\Gamma_{T}/\{\pm\mathrm{Id}\}$. The natural map $\Gamma_{T}\to\mathrm{Aut}(\mathbb{D}_{T})$ descends to $P\Gamma_{T}\hookrightarrow\mathrm{Aut}(\mathbb{D}_{T})$. We consider the orbifold $P\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$ instead of $\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$. For $\omega\in\mathbb{D}_{T}-\mathcal{H}_{T}$, the stabilizer of $\omega$ in $P\Gamma_{T}$ is given by
$\mathrm{Stab}_{\omega}=\{g|_{Q_{F_{0}}}\,\big{|}\,g\in\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}}),\,g\cdot\omega=\omega\}/\{\pm\mathrm{Id}\},$
where $\omega$ represents a positive complex line in $Q_{F_{0}}\otimes\mathbb{C}$. There exists a $Z(F)\in\mathbb{P}\mathcal{V}_{T}$ satisfying $\mathscr{P}_{T}(Z(F))=[\omega]$. By definition of $\mathscr{P}_{T}$, there exists a path $\gamma_{F}\subset\mathbb{P}\mathcal{V}_{T}$ connecting $Z(F)$ and $Z(F_{0})$ such that $\gamma_{F}^{*}H^{2,0}(X_{F})=\omega$. Let $\varphi$ denote $\gamma_{F}^{*}\colon H^{2}(X_{F},\mathbb{Z})\to H^{2}(X_{F_{0}},\mathbb{Z})$. The homomorphism $\varphi$ satisfies $\varphi(H_{F})=H_{F_{0}}$, $\varphi(\Delta_{F})=\Delta_{F_{0}}$. Then we can define a group homomorphism
$\mathrm{Aut}(X_{F},H_{F})\to\{g|_{Q_{F_{0}}}\,\big{|}\,g\in\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}},\omega)\},\quad f\mapsto(\varphi\circ f^{*}\circ\varphi^{-1})|_{Q_{F_{0}}},$
which is surjective. It induces a surjective homomorphism $\mathrm{Stab}_{Z(F)}\to\mathrm{Stab}_{\omega}$.
Proposition  6.5.
The following statements are equivalent to each other.
  1. (i)
    The period map $\mathscr{P}_{T}\colon\mathcal{M}_{T}\to P\Gamma_{T}\backslash(\mathbb{D}_{T}-\mathcal{H}_{T})$ preserves the orbifold structures of the GIT quotient and the arithmetic quotient.
  2. (ii)
    The natural restriction map $\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}})\to\Gamma_{T}$ is an isomorphism.
  3. (iii)
    The group $\mathrm{O}(P_{F_{0}},H_{F_{0}},\Delta_{F_{0}})$ acts faithfully on $A_{P_{F_{0}}}$.
Proof.
The period map $\mathscr{P}_{T}$ preserves the orbifold sturctures if and only if for any $\omega\in\mathbb{D}_{T}-\mathcal{H}_{T}$ and an associated $Z(F)\in\mathbb{P}\mathcal{V}_{T}$, $\mathrm{Stab}_{Z(F)}\cong\mathrm{Stab}_{\omega}$. By Corollary 2.14, $\iota_{F_{0}}$ acts as $-1$ on $Q_{F_{0}}$, this is equivalent to say that the map
$\mathrm{Aut}(X_{F},H_{F})\cong\mathrm{O}(\Lambda_{F},H_{F},\Delta_{F},H^{2,0}(X_{F}))\to\{g|_{Q_{F_{0}}}\,\big{|}\,g\in\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}},\omega)\}$
is an isomorphism. Denote by $G_{Z(F)}\coloneqq\mathrm{O}(\Lambda_{F},H_{F},\Delta_{F},H^{2,0}(X_{F}))$ and $G_{\omega}\coloneqq\{g|_{Q_{F_{0}}}\,\big{|}\,g\in\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}},\omega)\}$.
By definition, there is a surjective group homomorphism $p\colon\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}})\to\Gamma_{T}$ defined by restriction. The stabilizer $G_{Z(F)}$ can be naturally embedded into $\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}})$ by the conjugation of $\varphi$. Via this embedding, $G_{Z(F)}$ is isomorphic to $p^{-1}(G_{\omega})$. Hence $G_{Z(F)}\cong G_{\omega}$ is equivalent to $\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}})\cong\Gamma_{T}$. The isomorphism $\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}})\cong\Gamma_{T}$ is equivalent to the condition that the identity element (hence every element) in $\Gamma_{T}$ admits a unique lift in $\mathrm{O}(\Lambda_{F_{0}},H_{F_{0}},\Delta_{F_{0}})$, which means that $\mathrm{O}(P_{F_{0}},H_{F_{0}},\Delta_{F_{0}})$ acts faithfully on $A_{P_{F_{0}}}$. ∎
6.2. Nodal Singular Types
As an application, we study the nodal singular types, i.e. the singular types with only nodes.
Proposition  6.6.
The orbifold structures are preserved in all nodal singular types.
The proposition will be concluded from the following Lemma 6.7 and Lemma 6.9.
Lemma  6.7.
For every irreducible nodal singular type, the orbifold structure is preserved.
Proof.
Let $r_{1},\cdots,r_{m}$ be the base of $L=A_{1}^{\oplus m}$. We have $\mathrm{O}(P,H,\Delta)\cong\mathrm{Aut}(\Delta)\cong\mathfrak{S}_{m}$ and $A_{P}\cong\mathbb{Z}/2\times(\mathbb{Z}/2)^{m}$. Moreover, $\mathrm{O}(P,H,\Delta)$ acts on $A_{P}$ by permuting the $m$ generators of the later $m$ $\mathbb{Z}/2$-factors. This action is faithful. ∎
Lemma  6.8.
Fix an arbitrarily nodal singular type. Let $r_{1},\cdots,r_{m}$ be the base of $L=A_{1}^{\oplus m}$. For any $v\in P$ such that $v=\mu H+\sum\limits_{i}\lambda_{i}r_{i}$, either all $\lambda_{i}$ are integers, or there exist at least five of $\lambda_{i}$ belonging to ${1\over 2}\mathbb{Z}\backslash\mathbb{Z}$.
Proof.
For any irreducible nodal singular type, the statement is followed by $P=\langle H\rangle\oplus L$.
Consider a reducible nodal singular type. Denote by $C_{1},\cdots,C_{l}$ the irreducible components of a sextic curve of the given singular type. In reducible cases, the statement is equivalent to saying that for any $C_{i_{1}},\cdots,C_{i_{k}}$ with $1\leq k<l$, $C_{i_{1}}+\cdots+C_{i_{k}}$ has at least five coefficients relative to $r_{1},\cdots,r_{m}$ belonging to ${1\over 2}\mathbb{Z}\backslash\mathbb{Z}$.
Each $C_{i}$ can be expressed as
$C_{i}={1\over 2}(d_{i}H+\sum_{j\in I_{i}}r_{j}),$
where $d_{i}=\deg C_{i}$, $I_{i}$ corresponds to the set of nodes arising from the intersection of $C_{i}$ with the union of other irreducible components. It follows that $C_{i_{1}}+\cdots+C_{i_{k}}$ satisfies
$C_{i_{1}}+\cdots+C_{i_{k}}\equiv{1\over 2}(\sum_{n=1}^{k}d_{i_{n}})H+{1\over 2}\sum_{j\in I_{i_{1},\cdots,i_{k}}}r_{j}\mod\langle H\rangle\oplus L,$
where $I_{i_{1},\cdots,i_{k}}$ corresponds to the set of nodes arising from the intersection of $C_{i_{1}}\cup\cdots\cup C_{i_{k}}$ with the union of other irreducible components. By Bézout theorem, the cardinality of $I_{i_{1},\cdots,i_{k}}$ is at least five, thus the lemma follows. ∎
Lemma  6.9.
For every reducible nodal singular type, the orbifold structure is preserved.
Proof.
Let $g\in\mathrm{O}(P,H,\Delta)$ such that $g^{*}$ acts on $A_{P}$ trivially. By Proposition 6.5, it is enough to show $g=\mathrm{Id}$, namely, $g$ preserves all $r_{i},1\leq i\leq m$. For each $r_{i}$ arising from a self-intersection node of an irreducible component, $r_{i}$ is preserved for the same reason as in Lemma 6.7. Thus it is enough to show that each $r_{i}$ arising from an intersection of two irreducible components is preserved. Let $C,C^{\prime}$ be two different irreducible components. Denote by $J\coloneqq C\cap C^{\prime}$.
If $|J|=1$, then both of $C$ and $C^{\prime}$ must be two lines. Denote by $r$ the associated exceptional curve. Consider ${1\over 2}(H+r)\in P\otimes\mathbb{Q}$, which is actually an element in $P^{\vee}$. Since $g^{*}=\mathrm{Id}$ on $A_{P}$, we have $g^{*}[{1\over 2}(H+r)]=[{1\over 2}(H+r)]=[{1\over 2}(H+g\cdot r)]$ in $A_{P}$, then $[{1\over 2}(r-g\cdot r)]=0$ in $A_{P}$. If $g^{*}r\neq r$, then ${1\over 2}(r-g\cdot r)\neq 0$ lies in $P$, which contradicts to Lemma 6.8.
Consider the case $|J|\geq 2$. We first assume that for any $p_{i},p_{j}\in J$, the associated exceptional curves $r_{i},r_{j}$ not form a $g$-orbit in $P$. Consider ${1\over 2}(r_{i}+r_{j})\in P^{\vee}$. We have $g^{*}[{1\over 2}(r_{j}+r_{j})]=[{1\over 2}(r_{j}+r_{j})]=[{1\over 2}(g\cdot r_{i}+g\cdot r_{j})]$ in $A_{P}$. If $g\cdot r_{i}\neq r_{i}$, then ${1\over 2}(r_{i}+r_{j}-g\cdot r_{i}-g\cdot r_{j})\neq 0$ belongs to $P$, which contradicts to Lemma 6.8.
Otherwise, there exist $p_{i},p_{j}\in J$ such that $g\cdot r_{i}=r_{j},g\cdot r_{j}=r_{i}$. If $|J|\geq 3$, then there exists a $p_{k}\in J\backslash\{p_{i},p_{j}\}$. Consider ${1\over 2}(r_{i}+r_{k})\in P^{\vee}$. Similarly, we have ${1\over 2}(r_{i}-r_{j}+r_{k}-g\cdot r_{k})\neq 0$ belonging to $P$. Contradiction.
Otherwise, $|J|=2$, then $C,C^{\prime}$ must be a line and a conic. Then there exist only three possible cases: one conic plus four lines, two conics plus two lines and one cubic plus one conic plus one line. Note that the third case actually contains more than one singular types, since the cubic could be smooth or nodal. But this is irrelevant to our discussion.
We can assume that $C$ is a line and $C^{\prime}$ is a conic. Let $C\cap C^{\prime}=\{p_{i},p_{j}\}$ and let $p_{k}$ be a node arising from the intersection of $C^{\prime}$ with an irreducible component of odd degree (which always exists). It is direct to verify that ${1\over 2}(H+r_{i}+r_{k})\in P^{\vee}$. Then we conclude that ${1\over 2}(r_{i}-r_{j}+r_{k}-g^{*}r_{k})\neq 0$ belongs to $P$. Contradiction. ∎
Remark  6.10.
It is interesting to know whether the orbifold structures are preserved for other singular types. This question is still open.
7. Examples and Applications
The moduli and period map for nodal singular types have been well studied, see [YZ23]. In this section, we describe some further examples as applications.
7.1. A Quintic Curve plus a Line
We consider sextic curves that split into a smooth quintic curve and a line. Let $V$ denote a complex vector space of dimension $3$.
7.1.1. $5$ Nodes
Suppose there is no tangent point. Denote by $T=T_{0}$. This case is first studied by Laza [Laz09], and is related to the study of the minimal-elliptic surface singularity $N_{16}$. This case is also included in [YZ23].
In this case, $\mathcal{V}_{T_{0}}$ consists of pairs $(F_{5},F_{1})\in\mathrm{Sym}^{5}V^{*}\times V^{*}$ such that $Z(F_{5})$ and $Z(F_{1})$ have five different intersection points. It is direct $\dim\mathcal{M}_{T_{0}}=14$, which coincides with the expected dimension.
For any fixed $F\in\mathcal{V}_{T_{0}}$, we have $L_{F}\cong A_{1}^{\oplus 5}$. Denote $\Delta_{F}$ by $\{e_{1},\cdots,e_{5}\}$. Corollary 5.4 implies $P_{F}=P_{F}^{\iota}$. By Proposition 5.9, we have:
Proposition  7.1.
The lattice $P_{F}$ is generated by $\langle H\rangle\oplus L_{F}$ and ${H-e_{1}-\cdots-e_{5}\over 2}$.
Remark  7.2.
For a generic $F\in\mathcal{M}_{T_{0}}$, the K3 surface $X_{F}$ admits an elliptic fibration without sections (it admits a bi-section). This fibration has $1$ type $I_{0}^{*}$ fiber and $18$ nodal fibers. The lattice $P_{F}$ is isomorphic to $U(2)\oplus D_{4}$. The Picard lattice of the associated Jacobian elliptic K3 surface is $U\oplus D_{4}$, and the Mordell–Weil group is trivial. This example also appears in [Bel97, Table 1.1, No. 6] and [CM24, Table 2].
Recall that $Q_{F}\coloneqq P_{F}^{\perp}$, and $Q_{F}\cong U(2)\oplus U\oplus D_{4}\oplus E_{8}$. The associated period domain is $\mathbb{D}_{T_{0}}\coloneqq\mathbb{D}(Q_{F})$, which is a type IV domain of dimension $14$. The associated arithmetic group $\Gamma_{T_{0}}$ has the following expression
$\Gamma_{T_{0}}=\{g\in\mathrm{O}(Q_{F})\,|\,g\textup{ acts on }A_{Q_{F}}\textup{ through the }\mathfrak{S}_{5}\textup{-action on }A_{P_{F}}\},$
where $\mathfrak{S}_{5}$-action permutes $e_{1},\cdots,e_{5}$.
The hyperplane arrangement $\mathcal{H}_{T_{0}}^{*}$ is empty. Thus $\mathscr{P}_{T_{0}}$ extends to an isomorphism $\mathcal{M}_{T_{0}}^{\mathrm{ADE}}\cong\Gamma_{T_{0}}\backslash\mathbb{D}_{T_{0}}$, see [Laz09, Theorem 4.1] and [YZ23, Proposition 3.7] [^1] There is a small typo in [YZ23, Proposition 3.7]: ”$\Sigma_{T}$” should be revised as ”$\overline{\Sigma}_{T}$”, the closure of $\Sigma_{T}$ in the GIT compactification. This criterion holds for all ADE singular types.. The Looijenga compactification with respect to the empty hyperplane arrangement is exactly the Baily–Borel compactification. Hence the period map $\mathscr{P}_{T_{0}}$ extends to an isomorphism $\mathcal{M}_{T_{0}}^{\mathrm{sst}}\cong\overline{\Gamma_{T_{0}}\backslash\mathbb{D}_{T_{0}}}^{\mathrm{bb}}$, also see [Laz09, Theorem 4.2] and [YZ23, Proposition 4.3]. In this case, the GIT compactification is normal, i.e. $\mathcal{M}_{T_{0}}^{\mathrm{sst}}=\overline{\mathcal{M}}_{T_{0}}$.
By Proposition 6.6, the period map $\mathscr{P}_{T_{0}}$ induces an isomorphism of $\mathcal{M}_{T_{0}}$ and $P\Gamma_{T_{0}}\backslash(\mathbb{D}_{T_{0}}-\mathcal{H}_{T_{0}})$ as orbifolds.
7.1.2. $1$ Tacnodes plus $3$ Nodes
Suppose there is only one tacnode. Denote by $T=T_{1}$. Let $\mathcal{V}_{5}$ consists of $F_{5}\in\mathrm{Sym}^{5}V^{*}$ such that $Z(F_{5})$ is a smooth quintic curve. Denote by $\mathbb{P}\mathcal{V}_{5,1}\to\mathbb{P}\mathcal{V}_{5}$ the universal family. In this case, $\mathbb{P}\mathcal{V}_{T_{1}}$ is a Zariski open subset of $\mathbb{P}\mathcal{V}_{5,1}$.
For any fixed $F\in\mathcal{M}_{T_{1}}$, we have $L_{F}\cong A_{3}\oplus A_{1}^{\oplus 3}$. Denote by $\{\epsilon_{1},\epsilon_{2},\epsilon_{3}\}$ the effective base of $A_{3}$ and by $e_{1},e_{2},e_{3}$ the effective bases of the three copies of $A_{1}$. Then $\Delta_{F}$ is given by $\{\epsilon_{1},\epsilon_{2},\epsilon_{3};e_{1};e_{2};e_{3}\}$.
Lemma  7.3.
We have $L_{F}^{-\iota}=P_{F}^{-\iota}$.
Proof.
By definition, $L_{F}^{-\iota}$ is generated by $\epsilon_{1}-\epsilon_{3}$, where $(\epsilon_{1}-\epsilon_{3})^{2}=-4$. Since $({\epsilon_{1}-\epsilon_{3}\over 2})^{2}=-1$ and $P_{F}$ is even, $L_{F}^{-\iota}$ does not admit any further saturation into $P_{F}$, namely $L_{F}^{-\iota}=P_{F}^{-\iota}$. ∎
Recall that $M_{F}$ denotes the lattice generated by $\langle H_{F}\rangle\oplus L_{F}$ and $u\coloneqq{H-(2\epsilon_{2}+\epsilon_{1}+\epsilon_{3})-e_{1}-e_{2}-e_{3}\over 2}$.
Proposition  7.4.
We have $P_{F}=M_{F}$.
Proof.
It is direct to check that there exists a decomposition of $M_{F}$ as following
$M_{F}=\langle H_{F}-e_{2},H_{F}-e_{3}\rangle\oplus\langle e_{1},u,\epsilon_{2},\epsilon_{1},\epsilon_{3}\rangle\cong U(2)\oplus D_{5}.$
Thus we have $A_{M_{F}}\cong(\mathbb{Z}/2)^{2}\times\mathbb{Z}/4$. By Lemma 7.3 and Proposition 5.9, the quotient $P_{F}\over{M_{F}^{\iota}\oplus M_{F}^{-\iota}}$ is isomorphic to a subgroup of $A_{M_{F}^{\iota}}$ and also a subgroup of $A_{M_{F}^{-\iota}}$. Since $A_{M_{F}^{\iota}}\cong(\mathbb{Z}/2)^{4}$ and $A_{M_{F}^{-\iota}}\cong\mathbb{Z}/4$, $P_{F}\over{M_{F}^{\iota}\oplus M_{F}^{-\iota}}$ is a subgroup of $\mathbb{Z}/2$. Moreover, we have ${M_{F}\over{M_{F}^{\iota}\oplus M_{F}^{-\iota}}}\cong\mathbb{Z}/2$. By definition, $M_{F}\subset P_{F}$, thus $M_{F}=P_{F}$. ∎
Remark  7.5.
For a generic $F\in\mathcal{M}_{T_{1}}$, the K3 surface $X_{F}$ admits an elliptic fibration $\pi_{1}$ with sections. This fibration has a type $I_{0}^{*}$ fiber and $18$ nodal fibers. The Mordell–Weil group of $\pi_{1}$ is of rank $1$ and torsion-free.
Remark  7.6.
For a generic $F\in\mathcal{M}_{T_{1}}$, $X_{F}$ also admits an elliptic fibration $\pi_{2}$ without sections (it admits a bi-section). This fibration has a type $I_{1}^{*}$ fiber and $17$ nodal fibers. The Jacobian elliptic surface of $\pi_{2}$ has Picard lattice isomorphic to $U\oplus D_{5}$. The Mordell–Weil group is trivial.
We have $Q_{F}\cong U\oplus U(2)\oplus A_{3}\oplus E_{8}$. The period domain $\mathbb{D}_{T_{1}}$ is a type IV domain of dimension $13$. The arithmetic group $\Gamma_{T_{1}}$ can be expressed as
$\Gamma_{T_{1}}=\{g\in\mathrm{O}(Q_{F})\,|\,g\textup{ acts on }A_{Q_{F}}\textup{ through the }\mathfrak{S}_{2}\times\mathfrak{S}_{3}\textup{-action on }A_{P_{F}}\},$
where $\mathfrak{S}_{2}$-action (respectively, $\mathfrak{S}_{3}$-action) permutes $\epsilon_{1},\epsilon_{3}$ (respectively, $e_{1},e_{2},e_{3}$). By [YZ23, Proposition 3.7], the hyperplane arrangement $\mathcal{H}_{T_{1}}^{*}$ is empty. The period map $\mathscr{P}_{T_{1}}$ extends to an isomorphism $\mathcal{M}_{T_{1}}^{\mathrm{ADE}}\cong\Gamma_{T_{1}}\backslash\mathbb{D}_{T_{1}}$ and further an isomorphism ¯$\overline{\mathcal{M}}_{T_{1}}\cong\overline{\Gamma_{T_{1}}\backslash\mathbb{D}_{T_{1}}}^{\mathrm{bb}}$.
7.1.3. $2$ Tacnodes plus $1$ Nodes
Suppose there are two tangent points. Denote by $T=T_{2}$. Due to a classical result by Schubert [Sch79], a generic smooth degree $d$ plane curve has $d(d-2)(d^{2}-9)\over 2$ bitangent lines. Thus a generic smooth plane quintic curve has $120$ bitangent lines. Denote by $\mathcal{M}_{5}$ the moduli space of smooth quintic curves. The natural projection $\mathcal{M}_{T_{2}}\to\mathcal{M}_{5}$ is a finite morphism of degree $120$.
For any fixed $F\in\mathcal{M}_{T_{2}}$, we have $L_{F}\cong A_{3}^{\oplus 2}\oplus A_{1}$. Denote by $(\epsilon_{i})_{i}$ and $(\delta_{i})_{i}$ the effective bases of two copies of $A_{3}$ respectively, and by $e$ the effective base of $A_{1}$. Then $\Delta_{F}$ is given by $\{\epsilon_{1},\epsilon_{2},\epsilon_{3};\delta_{1},\delta_{2},\delta_{3};e\}$.
Lemma  7.7.
We have $L_{F}^{-\iota}=P_{F}^{-\iota}$.
Proof.
By definition, $L_{F}^{-\iota}$ is generated by $\epsilon_{1}-\epsilon_{3}$ and $\delta_{1}-\delta_{3}$. The only possible value for $(a,b)$ $(0\leq a,b\leq 3)$ that results in $a(\epsilon_{1}-\epsilon_{3})+b(\delta_{1}-\delta_{3})\over 4$ having an even self-intersection number is $(a,b)=(2,2)$. However, $(\epsilon_{1}-\epsilon_{3})+(\delta_{1}-\delta_{3})\over 2$ is a root that is orthogonal to $H_{F}$, so it doesn’t belong to $P_{F}$. Thus $L_{F}^{-\iota}$ doesn’t admit any further saturation into $P_{F}$, namely $L_{F}^{-\iota}=P_{F}^{-\iota}$. ∎
In this case, $M_{F}$ is generated by $\langle H_{F}\rangle\oplus L_{F}$ and $u\coloneqq{H_{F}-(2\epsilon_{2}+\epsilon_{1}+\epsilon_{3})-(2\delta_{2}+\delta_{1}+\delta_{3})-e\over 2}$.
Proposition  7.8.
We have $M_{F}\cong\langle-4\rangle\oplus U\oplus D_{5}$.
Proof.
It is easy to check $A_{M_{F}}\cong(\mathbb{Z}/4)^{2}$, which is generated by ${H\over 2}+{3\epsilon_{1}+2\epsilon_{2}+\epsilon_{3}\over 4}$ and ${e\over 2}+{3\delta_{1}+2\delta_{2}+\delta_{3}\over 4}$. By [CS99, Chapter 15, §7.4], we have $A_{M_{F}}\cong 4^{-2}_{2}$ (Conway–Sloane symbol). By [Nik79, Cor 1.13.3], there exists a unique $N$ (up to isomorphism) of signature $(1,7)$ such that $A_{N}\cong 4^{-2}_{2}$. It is direct to verify that the discriminant form of $\langle-4\rangle\oplus U\oplus D_{5}$ is isomorphic to $A_{N}\cong 4^{-2}_{2}$. ∎
Proposition  7.9.
We have $P_{F}=M_{F}$.
Proof.
It is clear that $A_{M_{F}^{\iota}}\cong(\mathbb{Z}/2)^{4}$. By Lemma 7.7, $A_{M_{F}^{-\iota}}\cong(\mathbb{Z}/4)^{2}$. By a parallel argument in the proof of Proposition 7.4, we can conclude $P_{F}=M_{F}$. ∎
Remark  7.10.
For a generic $F\in\mathcal{V}_{T}$, the K3 surface $X_{F}$ admits an elliptic fibration $\pi_{1}$ with sections. There exists a type $I_{1}^{*}$ fiber and $17$ nodal fibers. The Mordell–Weil group is of rank one and torsion-free.
We have $Q_{F}\cong\langle 4\rangle\oplus A_{3}\oplus U\oplus E_{8}$. The arithmetic group $\Gamma_{T_{2}}$ can be described as
$\Gamma_{T_{2}}=\{g\in\mathrm{O}(Q_{F})\,|\,g\textup{ acts on }A_{Q_{F}}\textup{ through the }(\mathfrak{S}_{2})^{2}\times\mathfrak{S}_{2}\textup{-action on }A_{P_{F}}\},$
where the first two $\mathfrak{S}_{2}$-actions permute $\epsilon_{1},\epsilon_{3}$ and $\delta_{1},\delta_{3}$ respectively, and the last $\mathfrak{S}_{2}$-action permutes the two sets $\{\epsilon_{i}\}$ and $\{\delta_{i}\}$. Analogously to the previous cases, the period map $\mathscr{P}_{T_{2}}$ extends first to an isomorphism $\mathcal{M}_{T_{2}}^{\mathrm{ADE}}\cong\Gamma_{T_{2}}\backslash\mathbb{D}_{T_{2}}$ and then further to ¯$\overline{\mathcal{M}}_{T_{2}}\cong\overline{\Gamma_{T_{2}}\backslash\mathbb{D}_{T_{2}}}^{\mathrm{bb}}$.
7.2. A Quartic Curve with Two Bitangent Lines
The nodal case of quartic curves with two lines is discussed in [GMGZ18]. We consider the singular type $T$ when the two lines are both bitangent lines. Each smooth quartic curve admits $28$ bitangent lines. Let $\mathcal{M}_{4}$ denote the moduli space of quartic curves. The natural projection $\mathcal{M}_{T}\to\mathcal{M}_{4}$ is a finite morphism of degree $\binom{28}{2}$.
For any fixed $F\in\mathcal{M}_{T}$, we have $L_{F}\cong A_{1}\oplus A_{3}^{\oplus 4}$. Denote by $\Delta\coloneqq\{e;\alpha_{1},\alpha_{2},\alpha_{3};\beta_{1},\beta_{2},\beta_{3};\\ \epsilon_{1},\epsilon_{2},\epsilon_{3};\delta_{1},\delta_{2},\delta_{3}\}$. In this case, $M_{F}$ is generated by $\langle H\rangle\oplus L$, $u\coloneqq{H-e-(2\alpha_{2}+\alpha_{1}+\alpha_{3})-(2\beta_{2}+\beta_{1}+\beta_{3})\over 2}$ and $v\coloneqq{H-e-(2\epsilon_{2}+\epsilon_{1}+\epsilon_{3})-(2\delta_{2}+\delta_{1}+\delta_{3})\over 2}$.
Lemma  7.11.
We have $M_{F}^{-\iota}=P_{F}^{-\iota}$.
Proof.
Let $w\coloneqq{a(\alpha_{1}-\alpha_{3})+b(\beta_{1}-\beta_{3})+c(\epsilon_{1}-\epsilon_{3})+d(\delta_{1}-\delta_{3})\over 4}$ be an element in $(A_{3}^{\oplus 4})^{\vee}$, where $0\leq a,b,c,d\leq 3$. If $[w]$ is an isotropic element in $A_{A_{3}^{\oplus 4}}$, then we have $8\,|\,a^{2}+b^{2}+c^{2}+d^{2}$. Then either $(a,b,c,d)$ has two entries equal to $2$ and two to 00, or $(a,b,c,d)$ is equal to $(2,2,2,2)$. In the former case, all the entries correspond to roots orthogonal to $H_{F}$, none of which lies in $P_{F}$. The element corresponding to $(2,2,2,2)$ is contained in $M_{F}^{-\iota}$. ∎
Question  7.12.
Do we have $P_{F}=M_{F}$?
Remark  7.13.
Since $A_{M_{F}^{\iota}}\cong(\mathbb{Z}/2)^{6}$ and $A_{M_{F}^{-\iota}}\cong(\mathbb{Z}/4)^{2}\times(\mathbb{Z}/2)^{2}$, there may exist a saturation $M_{F}^{\iota}\oplus M_{F}^{-\iota}\hookrightarrow P$ such that ${P\over M_{F}^{\iota}\oplus M_{F}^{-\iota}}\cong(\mathbb{Z}/2)^{4}$. Notice that ${M_{F}\over M_{F}^{\iota}\oplus M_{F}^{-\iota}}\cong(\mathbb{Z}/2)^{3}$. There may not be an easy way to decide whether $P_{F}=M_{F}$.
Remark  7.14.
For a generic $F\in\mathcal{M}_{T}$, $X_{F}$ admits an elliptic fibration $\pi_{1}$ with sections. There exists a type $I_{1}^{*}$ fiber, $2$ type $I_{4}$ fibers and $15$ nodal fibers.
Remark  7.15.
For a generic $F\in\mathcal{M}_{T}$, $X_{F}$ admits another elliptic fibration without sections (it admits a bi-section). There exist $2$ type $I_{2}^{*}$ fibers and $10$ nodal fibers.
The arithmetic group $\Gamma_{T}$ can be characterized as
$\Gamma_{T}=\{g\in\mathrm{O}(Q_{F})\,|\,g\textup{ acts on }A_{Q_{F}}\textup{ through the }(\mathfrak{S}_{2})^{4}\times\mathfrak{S}_{4}\textup{-action on }A_{P_{F}}\},$
where the four $\mathfrak{S}_{2}$-actions permutes $\alpha_{1},\alpha_{3}$; $\beta_{1},\beta_{3}$; $\epsilon_{1},\epsilon_{3}$ and $\delta_{1},\delta_{3}$ respectively, and the $\mathfrak{S}_{4}$-action permutes the four sets $\{\alpha_{i}\}$, $\{\beta_{i}\}$, $\{\epsilon_{i}\}$ and $\{\delta_{i}\}$. Following the same argument, $\mathscr{P}_{T}$ extends to an isomorphism $\mathcal{M}_{T}^{\mathrm{ADE}}\cong\Gamma_{T}\backslash\mathbb{D}_{T}$ and further to ¯$\overline{\mathcal{M}}_{T}\cong\overline{\Gamma_{T}\backslash\mathbb{D}_{T}}^{\mathrm{bb}}$.
The moduli space of quartic curves admits an occult period map to an arithmetic ball quotient dimension $6$, see [Kon11]. A smooth quartic curve is a Riemann surface of genus three. Its Jacobian is a principal polarized abelian variety of dimension three. This leads to a birational equivalence between the moduli space of quartic curves and an arithmetic quotient of the Siegel upper half-space $\mathcal{S}_{3}$ of degree $3$. It would be interesting to study more relations among the three arithmetic quotients of different types and their compactifications.
7.3. Zariski Pairs
These examples had been discussed in [YX12, §3]. We present a construction of them from a more geometric viewpoint.
A pair of reduced plane curves of the same degree, which have the same combinatorial type of singularities but are not equisingular deformation equivalent, is usually called Zariski pairs. Zariski showed that the space of irreducible sextic curves with six cusps as the only singularities form at least two irreducible components, which provided the first example, see [Zar31]. Degtyarev showed that they actually form exactly two irreducible components, see [Deg08, Theorem 5.3.2]. Both of the two irreducible components are smooth of the same expected dimension.
Denote by $T_{1},T_{2}$ the two corresponding singular types, where $T_{2}$ corresponds to sextic curves whose six cusps lie on a conic.
Let $C$ be a sextic of $T_{2}$-type. Denote by $D$ for the conic containing the six cusps $p_{1},\cdots,p_{6}$ of $C$. By Bézout theorem, $D\cap C=\{p_{1},\cdots,p_{6}\}$ and the intersection multiplicity of each $p_{i}$ is two. Let $f\colon S\to\mathbb{P}^{2}$ be the blowup of $\{p_{1},\cdots,p_{6}\}$ on $\mathbb{P}^{2}$, and denote by $E_{1},\cdots,E_{6}$ the exceptional curves. We still denote by $C$ (respectively, $D$) the strict transform of $C$ (respectively, $D$). Then $C$ and $D$ are disjoint in $S$. For each $1\leq i\leq 6$, we have $E_{i}^{2}=-1$ and $E_{i}\cdot D=1$.
Let $\pi\colon X\to S$ be the double cover branched over $C$. Then $\pi^{-1}(D)$ is a disjoint union of $D_{1}$ and $D_{2}$, where $D_{1},D_{2}$ are both isomorphic to $D$. Moreover, we have $\pi^{*}E_{i}=\widetilde{E}_{i}+\widetilde{F}_{i}$ and $\pi^{*}C=2\widetilde{C}$, where each $(\widetilde{E}_{i},\widetilde{F}_{i})$ forms a $A_{2}$-type Dynkin diagram, $\widetilde{C}$ is reduced and $\widetilde{E}_{i}\cdot\widetilde{C}=\widetilde{F}_{i}\cdot\widetilde{C}=1$. Let $H$ be the pullback of the hyperplane class on $\mathbb{P}^{2}$. We have $H^{2}=2$ and $H\cdot D_{i}=2$. It is direct to verify that
$D_{1}=H-\frac{1}{3}\sum_{i=1}^{6}(2\widetilde{E}_{i}+\widetilde{F}_{i}),\quad D_{2}=H-\frac{1}{3}\sum_{i=1}^{6}(\widetilde{E}_{i}+2\widetilde{F}_{i}).$
It is clear that neither $D_{1}$ nor $D_{2}$ lies in $\left\langle H\right\rangle\oplus L$.
Due to the classification by Zariski and Degtyarev, we have the following:
Proposition  7.16.
  1. (i)
    For the $T_{1}$-type, the generic Picard group $P_{T_{1}}$ of the associated K3 surfaces is generated by $H$ and $L$.
  2. (ii)
    For the $T_{2}$-type, the generic Picard group $P_{T_{2}}$ of the associated K3 surfaces is generated by $H$, $L$ and $D_{1}$.
The singular type $T_{2}$ provides an example for $P\neq M$.
We have $Q_{T_{1}}=P_{T_{1}}^{\perp}\cong\langle-2\rangle\oplus U(3)^{\oplus 2}\oplus A_{2}^{\oplus 2}$, $Q_{T_{2}}=P_{T_{2}}^{\perp}\cong\langle-2\rangle\oplus U\oplus U(3)\oplus A_{2}^{\oplus 2}$. Both of the period domains $\mathbb{D}_{T_{1}}$ and $\mathbb{D}_{T_{2}}$ are type IV domains of dimension $7$. The arithmetic group $\Gamma_{T_{1}}$ can be written in the form of
$\Gamma_{T_{1}}=\{g\in\mathrm{O}(Q_{T_{1}})\,|\,g\textup{ acts on }A_{Q_{T_{1}}}\textup{ through the }(\mathfrak{S}_{2})^{6}\times\mathfrak{S}_{6}\textup{-action on }A_{P_{T_{1}}}\}.$
Similar for the arithmetic group $\Gamma_{T_{2}}$,
$\Gamma_{T_{2}}=\{g\in\mathrm{O}(Q_{T_{2}})\,|\,g\textup{ acts on }A_{Q_{T_{2}}}\textup{ through the }(\mathfrak{S}_{2})^{4}\times\mathfrak{S}_{4}\textup{-action on }A_{P_{T_{2}}}\}.$
For $i=1,2$, the period map $\mathscr{P}_{T_{i}}$ extends to an isomorphism $\mathcal{M}_{T_{i}}^{\mathrm{ADE}}\cong\Gamma_{T_{i}}\backslash\mathbb{D}_{T_{i}}$ and further to ¯$\overline{\mathcal{M}}_{T_{i}}\cong\overline{\Gamma_{T_{i}}\backslash\mathbb{D}_{T_{i}}}^{\mathrm{bb}}$.